A Bibliography of Publications in International Journal of Computer Processing of Oriental Languages and International Journal of Computer Processing of

Languages

Nelson H. F. Beebe University of Utah

Department of Mathematics, 110 LCB

155 S 1400 E RM 233 Salt Lake City, UT 84112-0090

USA

Tel: +1 801 581 5254 FAX: +1 801 581 4148

E-mail: beebe@math.utah.edu, beebe@acm.org,
beebe@computer.org (Internet)
WWW URL: http://www.math.utah.edu/~beebe/

14 October 2017 Version 2.08

Title word cross-reference

3 [ZMWP04]. k [Kwo03, Sar09]. N [TM00, KNL06].

-gram [TM00]. -grams [KNL06]. -means [Sar09]. -Nearest [Kwo03]. -NN [Sar09].

2002 [CD03]. 2007 [Lu08]. 22nd [Li09].

5th [Lee04].

AAC [BC06b, Hil06, LS06]. Abbreviations [FLW07]. Abductive [BL02]. Abnormality [KLL04]. AC [BC06a]. Access [LP02, Sak05]. Accommodating [SL03]. Accuracy [NH05]. ACE [LQLY07]. Acoustic [AMLJ03]. Acoustics [RC05]. Acquisition [CLLC08]. across [PLYZ95]. Act [LS02b]. Ad [KASAZS04]. Ad-hoc [KASAZS04]. Adaptive [ATAAAS11, Lee02, Wak02]. Address [WHSN04]. Advanced [Sak05, Won05, iZL07]. Affine [Wak02]. Agent [DSL^{+}02]. Agreement [ATWD09]. Alert [ASARAE12]. Algorithm [KJLL09, MH00, WCC01]. Algorithmically [AA04b]. Algorithms [GW06, LKSL08, Sar09]. alignment [KY09]. Alternative [BC06b, Hil06, Yon06]. Ambiguity [HLBL00, KZCJLL03, LHM^{+}00]. Analysis [AK09, KJLL09, LS02b, LRKL05, TKWL05, XK11b, YLC08]. Analyzer [TM00]. ANCS [AA04a]. Annotation [KLC^{+}00, LAM08, ZWT09]. Answer [KKSL01]. Answering [KKSL01, MKI07]. Application [FI00, Lee02, PWK02]. applications [CPJ93]. Appraisal [ZQCL11]. Approach [ARASGAS11, EB08, FLW07, HY07, Had07, HSBAS12, HAL09, KL04, Kwo03, Lee02, LWZF11, MBBS05, RZWX08, SKW^{+}00, SZSY07, WWCC07, YLC08, YCW11, ZMZ12]. Approaches [HKAH11, KC01, KNL06, ZZZL11, ZLZ11]. Appropriate [And06]. ARabic [BM08, AK09, AS04, ASC06, AS09a, AA04a, AA04b, ARASGAS11, ASARAE12, ATWD09, ATAAAS11, AZ04, CBA07, DAZF05, HY07, Had07, HH09, HASE11, IA09, KASAZS04, MC09, MZA11, MZA12, ONAK12, Sar09, SS07, SRMB04, STK07, ZMZ12]. Archives [TH12]. art [SS07]. Articles [Fai06]. Articulation [AMLJ03]. Asia [Ano04a, FZ03]. Asian [LL01]. Assessing [KY09]. Assisted [LTT02]. Association [ARASGAS11]. Attributes [CLLC08]. Augmentative [And06, BC06a, BC06b, Hil06, Yon06]. Author [Ano00, Ano01, Ano02, Ano03, Ano04b, Ano05, Ano06, Ano07, Ano08, Ano09, Ano11]. Automatic [FLW07, KCL03, KY09, OC03, STK07, SBN^{+}05, Tsu02]. Automatically [JLP04, LHM^{+}00]. Automatically-Extracted [LHM^{+}00].

BANGANUBAD [Sah05]. Bangla [Sah05]. Bank [XLWL09]. Based [ASC06, ASARAE12, BM08, CLTL02, CC02, CZZ^{+}05, CYL07, CJL12, DC09, FI00, GAH08, HASE11, HC02, HCR03, IKK^{+}09, JT00, KLC^{+}00, KKSL01, KL04, Kwo03, LW00, LS02a, LJL11, LKL01, LTZ^{+}01, LWLW06, MC09, MKI07, NIA06, ONAK12, SKW^{+}00, SSH05, SW00, ŠHHH08, SJM05, SCH11, TCC02, TT00, WLJ03, WHSN04, WL05, WXX^{+}08, XWW08, YCW11, ZZ11, ZMWP04, Ano04a, Cha01, CK02, FZ03, Hil06, HC12, IFK^{+}02, ITM^{+}07, Ish02, LWZF11, LH08, PWK02, SL03, SZSY07, WC01, WWCC07, WCC01]. Behavioural [WXL06]. Best [CD03, Lee04, Lu08]. between [ATWD09, KASAZS04, ONAK12]. Bi [CC01]. Bi-level [CC01]. Bigram [WL05]. Bilingual [KJLL09, Kwo01, ZMI05]. bitmap [PLYZ95]. Blended [LZ03]. Block [HC02]. Boolean [CLWK01]. Boundaries [CSC06]. Browsing [WCC00]. Buddhist [TH12]. Building [AKAAS11, BS11, CJL12, XLWL09, ZZLZ07, GAH08]. Business [WLR01].

Calculation [ZZ11]. CALL

[MZA11, MZA12]. Calligraphic [WLW03]. Calligraphy [CC01, HC01, HC02]. Caption [SKW^{+}00]. Case [HL12, WH11]. Categorial [JLP04, LP02]. Categorization [ARASGAS11, CZZ^{+}05, Kwo03]. Caused [KC01]. Cf [RL06]. Cf-ranking [RL06]. chī [WH11]. Chains [BL02]. Challenge [Hil06]. Channel [SW00]. Channel-Weighting [SW00]. Character [CLTL02, CC02, FKL02, IKK^{+}09, Kuo11, LAM08, Sar09, SSH05, TT00, VNL02, Wak02, WHSN04, HW92, LP96, Lun93, PLYZ95]. Characters [HG83, LKCK00, PHO02]. Chat [WXL06]. Checker [ATWD09, KKC04]. Checking [ATWD09]. China [Lee04]. Chinese [KL08, Lu08, And06, BC06a, CLTL02, CC02, CJL12, CAM08, CHC08, CLLC08, DC09, FLW07, GW06, GAH08, GW12, HG83, Hsi11, HW92, HC01, HC02, JT00, JNKL08, KZCJLL03, Kuo11, Kwo01, LKCK00, Lee02, LWY00, LWY03, LRKL05, LQLY07, LTB08, LYY01, LC12, LS06, LJAL02, LAM08, MW01, MWW01, SKW^{+}00, SSH05, ŠHHH08, TKWL05, TCC02, TT00, WCC00, WT02, WHSN04, WL05, WWCC07, WCC01, WLR01, WLYZ02, WXL06, XLWL09, Yon06, YCWL01, ZN04, ZMI05]. Chinese-to-Korean [KZCJLL03]. Chunk [WWCC07]. Chunking [DC09, HCR03]. Chunks [KK05]. Citizen [Jin00]. Classification [CAM08, Kim06a, KY09, Kwo03, LAM08, WLYZ02, XWW08, ZHW08, ZZZL11, ZQCL11]. Classifier [AA04a, LP96]. Classifiers [HL12]. Clause [Kim06b, KL04]. Clauses [Kim06a]. Clinical [Yon06]. CLSW2007 [Lu08]. Clustering [SJM05]. Co [LHM^{+}00, Yoo02]. Co-Occurrence [Yoo02, LHM^{+}00]. Code [JT00]. Coding [HC02, IA09, LJAL02, TT00]. Coercion [HHA08]. Cognition [LH08]. Cohesion [MKI07]. Collocation [GW12, XLWL09, ZZLZ07]. Color [SKW^{+}00]. Combination [KNL06, Sar09]. Combinatory [JLP04, LP02]. Combined [WWCC07]. Combining [LYK03]. Communication [And06, BC06a, BC06b, Hil06, Yon06]. Communities [Jin00]. Comparative [ZZZL11]. Comparison [AS04, KASAZS04, ONAK12, Yon06]. Complex [ITM^{+}07, IIK^{+}11]. Components [LWY00]. Composite [KCL03]. Compositional [ITM^{+}07]. Compositionality [Had07]. Compound [ITM^{+}07, MH00]. Comprehensive [ZMWP04]. Compression [CC01, TT00]. Computational [HSBAS12]. Computer [LD04, Li09, LTT02]. Computer-Assisted [LTT02]. Computing [SM02]. Concatenative [MBBS05]. Concealment [HC02]. Concept [LHM^{+}00]. Conceptual [GAH08, LH08, MMFA00]. Conditional [ZHW08]. Conference [Li09]. Conflation [AS04]. Conflict [JT00]. Constraints [YYW01]. Construct [CHY03]. Constructing [JLP04]. Construction [QLZ^{+}08, ZMI05]. Constructions [IIK^{+}11]. Content [SKW^{+}00, TKWL05, WC01]. Content-Based [SKW^{+}00, WC01]. Context [EB09, HLY^{+}01, IFK^{+}02]. Context-free [IFK^{+}02]. Contextual [MKI07, YLC08]. Continuous [BS11, PTCC05]. Contour [TT00]. Contrastive [LRKL05]. Conveying [KJLL09]. Cooperative [LD04]. CORBA [WLJ03]. Core [LS06]. Coreference [RZWX08]. Corpora [JNKL08]. Corpus [GAH08, HKAH11, HC12, KLC^{+}00, LDZ11, PTCC05, QLZ^{+}08, Yoo02, ZLZ11, AS09b, BS11]. Corpus-Based [GAH08, HC12]. Corpus-driven [HKAH11]. Correcting [MC09]. Correction [HY07, TM00]. Correspondence [SSH05]. Counted [ATWD09]. Court [TLSC06]. Coverage [TKWL05]. Crawled [KOK03]. Cross [ASC06, FI00, KNL05, Kwo01, NIA06, Sak01]. Cross-Cultural [NIA06]. Cross-Language [ASC06, FI00, KNL05, Sak01]. CSD [ZZ11]. CSRA [WWCC07]. Cube [Hsi11]. Cultural [LZ03, NIA06].

D [ZMWP04]. Data [CHY03, FK03, LS06, PWK02, Yoo02, ZWT09]. Database [AKAAS11, LP02, MW01]. Databases [LKCK00, MWW01]. DEAR [BM08]. DEAR-ONTO [BM08]. Decision [HCR03, LS02b]. Decomposition [HCR03, NH05, PHO02, SW00]. Dependency [Kim06b, LJL11, SNT09]. depth [BL02]. DErivational [BM08]. Design [CLTL02, HG83, LWY03, LTT02]. Designing [LKCK00]. Detecting [ASARAE12]. Detection [HY07, HZLL09, LQLY07]. Determination [HLY^{+}01]. Determine [Kim06b]. Deterministic [Yoo02]. Developing [And06, LS06]. Development [Hil06, ITM^{+}07]. Developments [SL03]. Dialogue [Cha01]. Dictation [IKK^{+}09]. Dictionary [CZZ^{+}05, ITM^{+}07, KJLL09, ONAK12, YCW11, ZMI05]. Dictionary-Based [CZZ^{+}05]. Diffusion [WLW03]. Digital [TH12]. Digitization [CBA07]. Dimension [LC12]. Disambiguation [Lee02, LYK03, ZMZ12]. Discontinuous [ML02]. Discussion [LD04]. Distance [CCLC04, HCK^{+}02, Li00]. Distributed [YYW01]. Distribution [KY09, LWLW06]. Diversity [SL03]. Document [CC02, Ish02, KNL06, LTZ^{+}01, TT00, WC01]. Documents [ASC06, CBA07, JLP04, KASAZS04, SBN^{+}05, SS01, ZN04]. Domain [AZ04, CZZ^{+}05, GAH08]. driven [HKAH11, WCC01]. DSK [Oka07]. Dynamic [LJL11].

e-Learning [FK03]. E-Mail [ASARAE12]. Easily [Kim06a]. Editor [AS09a, Lee04, Li08, Li09, Lu08, Won05, Zho11]. Editors [Ano04a, BC06a, CD03, FZ03, FKL02, HH12, KL08, LL01, MW01, SM00, SM02, XK11a, YY11, iZL07]. Education [HCK^{+}02]. Educational [Ano04a, FZ03]. Effect [CAM08]. Effective [FK03, RZWX08, YCWL01]. Effectuation [BNS11]. Efficient [LWZF11, MH00, Yoo02]. Effort [TLSC06]. Electronic [KSH00]. Ellipsis [HH09]. Empirical [ZZZL11]. Enabled [ZMWP04]. Encoding [CLTL02, Lun93]. Engine [CLTL02, CC02, PWK02, ZN04]. Engineering [TLSC06]. English [ASC06, GW06, JNKL08, KC01, Kwo01, Sah05, Sak01, SRMB04, Tsu02, Yon06, ZMI05]. Enhanced [WHSN04]. Enhancement [NH05]. Enterprise [YYW01]. Entity [ENB07, EB08, LQLY07, ŠHHH08]. Entropy [CYL07, EB08, ZHW08]. Environment [Jin00, WLJ03]. Environments [SL03]. Error [HC02, TM00, WCC01]. Error-driven [WCC01]. Errors [Ish02, MC09]. Evaluating [ASARAE12, AS09b]. Evaluation [KY09, Oka07]. Event [CAM08, Fai06, HHA08, HZLL09, HAH12, LWY03, LWLW06, LDZ11]. Event-Based [LWLW06]. Events [AKAAS11, WT02]. Every [Jin00]. Every-Citizen [Jin00]. Evidence [Hil06]. Evidence-based [Hil06]. Evidentiality [SCH11]. Examining [RL06]. Example [CYL07]. Example-Based [CYL07]. Examples [Wan04, ZWT09]. Expansion [FLW07]. Experience [LKCK00, LQLY07]. Experimental [GW06]. Expert [LTT02]. Exploitation [DAZF05]. Exploiting [Kwo01]. Exploring [SS11]. Expression [LRKL05, ZQCL11]. Expressions [MMFA00, SZSY07]. Extendable [MH00]. Extracted [LHM^{+}00]. Extraction [AK09, Fai06, Ish02, JNKL08, LW00, LWY03, Li08, OC03, PHO02, SS01, Tsu02, WHSN04, WCC01, WXLY05].

Fast [KKSL01]. FCA [CLLC08]. Feature [AK09, CJL12, LRKL05, SZSY07]. Feature-based [SZSY07]. Features [CC02, Kim06a, LJL11, XK11b]. Feedback [Sak01]. file [PLYZ95]. Filtering [KASAZS04, LTZ^{+}01]. Filters [ASARAE12]. Financial [LWY03]. Font [CHC08, SSH05, PLYZ95]. Foreign [KC01, OC03]. formats [PLYZ95]. Frame [LH08]. Frame-based [LH08]. Framework [JLP04, LD04, TH12]. free [IFK^{+}02]. Frequency [LS06]. Future [SL03].

Generalized [Had07]. Generated [JLP04]. Generation [KCL03, LKL01, STK07, SSH05]. Global [Wak02]. gram [AS04, TM00]. Grammar [IFK^{+}02, JLP04, LP02, YCWL01]. Grammar-based [IFK^{+}02]. grams [KNL06]. Granularity [SS11]. Graph [LJL11]. Graph-Based [LJL11].

Graphemes [CK02]. Guest [AS09a, Ano04a, BC06a, CD03, FZ03, FKL02, HH12, KL08, LL01, Lee04, Li08, Li09, Lu08, MW01, SM00, SM02, Won05, XK11a, YY11, iZL07, Zho11].

Han [PLYZ95]. Handwriting [SS07]. Handwritten [AK09, CK02, HAL09, PHO02, SSH05, Wak02, WHSN04]. Hangul [CK02]. Hashing [MH00]. HCI [FK03]. HELP [GW12]. Hidden [LS02b]. Hierarchical [Kwo03, ZZZL11, LP96]. Hindi [HAL09]. Hints [AA04b]. Historical [CBA07]. history [Lun93]. HMM [CK02]. HMM-based [CK02]. hoc [KASAZS04]. Holistic [WHSN04]. Homophone [Lee02]. Homophonic [CLTL02]. Hong [TLSC06]. HTML [ZN04]. Huffman [TT00]. Human [NH05]. Human-Written [NH05]. Hybrid [CLWK01, FLW07, HY07, HASE11, KLL04, Kim06b, ZMZ12].

ICCLC [CD03]. ICCPOL [Li09]. Icon [And06]. Identification [CAM08, CSC06, HASE11, LTB08, SBN^{+}05, WL05]. Identifying [Fai06, HSBAS12, LWY00]. II [Ano04a]. Image [CC02, HC01, SKW^{+}00, SBN^{+}05, VNL02]. Images [Ish02, TT00]. Imperatives [SL03]. Implementation [LTT02]. Improved [IA09, MC09]. Improvement [MBBS05]. Improves [EB09]. Improving [CZZ^{+}05, LJAL02, ZLZ11]. In-Vocabulary [WHZ08]. incompatible [PLYZ95]. Increasing [SMB04]. Incremental [KOK03, LKSL08]. Index [Ano00, Ano01, Ano02, Ano03, Ano04b, Ano05, Ano06, Ano07, Ano08, Ano09, Ano11]. Indexing [CLWK01, KKSL01, STK07, YCW11]. Indian [EB08, MBBS05]. Individualized [Wan04]. Inductive [Wan04]. Inferring [LDZ11]. Influence [RC05]. Information [ASC06, AKAAS11, CTL00, Fai06, FI00, Ish02, JLP04, KC01, KNL05, KNL06, Kwo01, LL01, LWY00, LHM^{+}00, LD04, LWLW06, Sak01, Sak05, SJM05, SZSY07, SS01, TT00, Won05, WXLY05, YYW01, YLC08, YCW11, ZN04]. Ingestion [HHA08]. Inheritance [CC01]. Ink [WLW03]. Input [JT00, Kuo11, TCC02]. Inquiry [Wan04]. Integrated [TH12]. Integrating [KLC^{+}00, LWLW06, YYW01]. Intellectual [TCC02]. Intelligent [CBA07, CHY03, Kuo11]. Inter [Kim06b]. Inter-Clause [Kim06b]. Interactive [HCK^{+}02]. Interactivity [FK03]. Interface [MWW01]. Intermediary [ZMI05]. International [Li09]. Internet [FI00, Li00]. Internet-Mediated [Li00]. Interpretation [LP02, NIA06]. Intonation [JLP04]. Introduction [AS09a, Ano04a, BC06a, CD03, FZ03, FKL02, HH12, KL08, LL01, Lee04, Li08, Li09, Lu08, MW01, SM00, SM02, Won05, XK11a, YY11, iZL07, Zho11]. Investigation [CTL00, WCC01]. Issue [AS09a, Ano04a, BC06a, CD03, FZ03, FKL02, KL08, LL01, Lee04, Li08, Li09, Lu08, MW01, SM00, SM02, Won05, iZL07]. Issues [LZ03].

Japanese

[ITM^{+}07, IKK^{+}09, IIK^{+}11, Lun93, ML02, Sak01, TM00, Tsu02, Wak02, ZMI05]. Japanese-Chinese [ZMI05]. Japanese-English [Sak01]. Japanese-Katakana [Tsu02]. jìn} [HC12]. Joint [Lee04]. Jurilinguistic [TLSC06].

Kana [IKK^{+}09]. Kanji [IKK^{+}09].

Katakana [Tsu02]. Key [Kuo11, WHSN04]. Keyboards [Oka07]. Keyword [FI00]. Knowledge [KLC^{+}00, MKI07]. Kong [TLSC06]. Korea [Lee04]. Korean [HLBL00, KC01, KNL06, KCL03, KZCJLL03, KKC04, LS02b, LJL11, LRKL05, ML02, OC03, PHO02, RL06, Yoo02]. Korean-to-Japanese [ML02].

Labeling [DC09]. Labels [KCL03]. Laboratory [DSL^{+}02, WLJ03, CCLC04]. Language [ASC06, AS09a, ATWD09, ATAAAS11, BS11, BC06a, BC06b, FI00, HH09, IFK^{+}02, ITM^{+}07, KNL05, KJLL09, Kwo01, KL08, LP02, Lee04, LKL01, LTZ^{+}01, MC09, MWW01, Sak01, SNT09, WC01, WXL06, iZL07]. Languages [EB08, LL01, Li09, MBBS05, RC05, SM02]. Large [DAZF05, LDZ11, ZZZL11, LP96]. Large-Scale [ZZZL11]. large-set [LP96]. Law [BNS11]. LDC [Kwo01]. Learner [SL03]. Learning [CCLC04, EB09, FK03, Jin00, KL04, LKCK00, LZ03, Lee02, Li00, LKSL08, LTT02, SL03, SMB04, WLW03, Wan04, WCC01, XK11b, YLC08, YCWL01, ZWT09, ZMWP04]. Letters [AK09]. Level [AZ04, SS11, XK11b, CC01, LKL01]. Lexical [Hsi11, LHM^{+}00, Lu08, Yoo02]. Lexicalized [TH12]. Lexico [KKSL01]. Lexico-Syntactic [KKSL01]. Lexicon [KJLL09]. Light [IIK^{+}11]. Like [LW00, Li00]. Likelihood [LW00]. Line [LYY01, CK02, SS07, VNL02]. lingual [Cha01, LP96]. Linguistic [HKAH11, KLC^{+}00, WXL06]. Link [KJLL09, LH08]. Live [LKSL08]. Local [WL05]. Locating [WL11]. Location [CC01, LDZ11]. Logical [Had07]. Lossless [CC01]. Low [JT00]. Lyric [XWW08]. Lyric-Based [XWW08].

Machine [CYL07, KZCJLL03, KLL04, KL04, KY09, LRKL05, ML02, NSNN07, Sak01, SRMB04, Wan04, YLC08, ZLZ11]. Machine-Learning [Wan04]. Macro [CAL^{+}00, DSL^{+}02]. Mail [ASARAE12]. Man [LD04]. Man-Computer [LD04]. Mandarin [And06, BC06a, BC06b, GW12, HL12, HC12, LTB08, LS06, LH08, SW00, WCC00, Yon06]. Many [HL12]. Map [WL11]. Mapping [GAH08, LTB08]. Market [WXX^{+}08]. Markov [LS02b]. MARVS [HHA08, HAH12, WH11]. Matching [AS04, KKSL01, KLL04, WHSN04]. Matter [LS02a]. Maximal [LJAL02]. Maximum [CYL07, EB08, ZHW08]. means [Sar09]. Measures [LYK03]. Mediated [Li00]. Message [Kuo11]. Messages [ASARAE12]. Metafont [HG83]. Metaphors [GAH08]. Method [Ish02, JT00, Kim06b, SW00, YCWL01, ZZ11]. Methods [RL06]. Mid [BL02]. Mid-depth [BL02]. Mining [ARASGAS11, CHY03, Li08, LKSL08, SZSY07, WT02, WXX^{+}08]. Mismatch [KC01]. Mobile [DSL^{+}02, KSH00, PWK02, TCC02]. Modal [LRKL05]. Model [CZZ^{+}05, CYL07, CSC06, HCK^{+}02, HZLL09, HCR03, IFK^{+}02, ITM^{+}07, Ish02, Kuo11, LS02b, NSNN07, OC03, QLZ^{+}08, SNT09, TM00, WL05, WLR01, XWW08]. Model-based [Ish02]. Modeling [FK03, ZHW08]. Modelling [AMLJ03]. Models [BS11, CLWK01, MC09, WC01]. Morpheme [KLC^{+}00]. Morphological [DAZF05, HLBL00, STK07, TM00]. Morphology [AA04a]. MSA [AS09b]. MT [LKL01]. Multi [Cha01, LP96, LKL01, ML02]. Multi-level [LKL01]. Multi-lingual [Cha01, LP96]. Multi-Word [ML02]. Multiengine [EB09]. Multilingual [Jin00]. Multimedia [SM02]. Multiple [LYK03, XK11b]. Music [YCW11]. Mutual [SZSY07].

N [AS04]. N-gram [AS04]. Name [GW06]. Named [ENB07, EB08, ŠHHH08]. Names [AA04b, AZ04, ONAK12]. Natural [AS09a, LP02, LKL01, LTZ^{+}01, MWW01, SNT09]. Near [GW12]. Near-synonymy [GW12]. Nearest [Kwo03]. Neighbor [Kwo03]. NER [EB09]. Net [ZZLZ07]. Network [KK05, Lee02, HW92, LP96]. Networks [CK02, CPJ93]. Neural [CPJ93, CK02, Lee02, LP96]. Newly [KOK03]. News [Fai06, FLW07, LWY03, TKWL05, WT02]. Nine [AMLJ03]. NLP [MZA11, MZA12]. NN [Sar09]. Non [AA04b, HY07, ITM^{+}07]. Non-Compositional [ITM^{+}07]. Non-Standards [AA04b]. Non-Words [HY07]. Normalization [IIK^{+}11, Wak02]. Notebook [KSH00]. Noun [AA04a, CSC06, SRMB04, WCC01]. Novel [HC01, LJAL02]. Numerals [ATWD09].

Objects [ATWD09, SMB04]. Obtainable [Kim06a]. Occurrence [Yoo02, LHM^{+}00]. OCR [Ish02, MC09, TM00]. Off [LYY01, SS07, VNL02]. Off-Line [LYY01, SS07, VNL02]. On-line [CK02, VNL02]. One [KJLL09, Kuo11, HL12]. One-Character-One-Key [Kuo11]. Online [CCLC04, LZ03, LD04]. ONTO [BM08]. Ontology [BM08, Cha01, CLLC08, TH12]. Ontology-based [Cha01]. Opinion [Li08, LWZF11, SZSY07, WXX^{+}08, XK11b]. Oriental [FKL02, Lee04, Li09, RC05, SM02, SBN^{+}05, Won05, LP96]. Other [SBN^{+}05, LC12]. Outcomes [Yon06]. Overview [MWW01, WXLY05].

Pacific [Ano04a, FZ03]. PageRanking [KOK03]. Pages [KOK03, WCC00]. Pair [JNKL08]. Pairs [Tsu02]. pào} [HC12]. Papers [CD03, CL02, KCL02, KL08, Lee04, Li09, Lu08]. Parallel [ASC06, CSC06, JNKL08]. Paraphrasing [IIK^{+}11]. Parser [ATAAAS11]. Parsing [HLBL00, KCL03, LJL11, SNT09, Yoo02]. Part [Ano04a, BNS11, FZ03, KCL03]. Part-of-Speech [KCL03]. Partition [CZZ^{+}05]. Partitioning [TT00]. Passage [LS02a]. Path [ZQCL11]. Pattern [EB09, ITM^{+}07, KKSL01, Lee04, WXX^{+}08]. Pattern-Based [WXX^{+}08]. Patterns [TT00, VNL02]. Performance [EB09, Oka07]. Periodic [WT02]. Persian [BS11]. Personalizing [CHC08].

Perspective [HHA08, HAH12, Yon06].

Peykare [BS11]. Phoneme

[GW06, JT00, KK05]. Phoneme-Based [JT00]. Phones [TCC02]. Phonetic [TCC02]. Phrase [LW00, WCC01]. Phrase-Like-Unit-Based [LW00]. Phrases [CSC06, SRMB04]. Picture [NIA06]. Picture-Based [NIA06]. Platform [CCLC04]. Point [AS09b]. Polysemy [WH11]. Possible [LZ03]. Post [EB09]. Post-processing [EB09]. Practice [Hil06]. Predicates [IIK^{+}11]. Prediction [HKAH11, WXX^{+}08, HW92]. Predictive [KKSL01]. Prefix [LJAL02]. Prevention [HC02]. Principles [GAH08]. Probabilistic [CSC06, HCR03, TM00]. Probabilities [KNL05]. Problem [KC01]. Processes [WLR01]. Processing [AS09a, ATAAAS11, CBA07, Lee04, LKL01, Li09, Won05, EB09, KL08]. Producing [AA04b]. Profile [HL12]. Programming [CCLC04]. Progressive [HC01]. Pronominalization [RL06]. Propagation [CC01]. Proper [MBBS05]. Prosody [GW12]. Pseudo [Sak01]. Pseudo-Relevance [Sak01].

Qualitative [SKW^{+}00]. Quality

[MBBS05, ZLZ11]. Quantifiers [Had07]. Quantitative [MMFA00]. Queries [LP02]. Query [FI00, KNL06, ZN04]. Question [KKSL01]. Question-Answering [KKSL01]. Questions [MKI07]. Quranic [HSBAS12]. Qwerty [Oka07].

ranking [RL06]. Rate [JT00]. Ratio [LW00]. Real [HCK^{+}02]. Real-time [HCK^{+}02]. Realistic [WLW03]. Recognition [BNS11, BS11, CK02, DAZF05, EB08, FKL02, IFK^{+}02, PTCC05, RC05, Sar09, SW00, ŠHHH08, VNL02, Wak02, WHSN04, ZQCL11, LP96, Lee04]. Recordings [STK07]. Rectangular [TT00]. recurrent [HW92]. Region [TT00].

Regular [CL02, KCL02]. Relation [CAM08]. Relational [LP02]. Relative [Kim06a]. Relevance [Sak01]. Relevant [Fai06]. Reliability [SCH11]. Reliable [KKSL01]. Rendering [PWK02]. Representation [Had07, Hsi11, WH11, WLYZ02]. Representations [MMFA00]. Representative [ZWT09]. Requisite [BNS11]. Research [Hil06, LTB08, SCH11]. Resolution [HH09, HLBL00, KC01, LHM^{+}00, RZWX08]. Resolving [KZCJLL03]. Resource [CJL12]. Retrieval [ASC06, AKAAS11, CC02, FI00, KASAZS04, KC01, KNL05, KNL06, Kwo01, LL01, LD04, Li08, LWZF11, LJAL02, MH00, Sak01, SKW^{+}00, WC01, YCW11]. Retrieving [ZN04]. Retrospective [WT02]. Rhetorical [SJM05]. Rim [Ano04a, FZ03]. Role [DC09]. Romanization [AA04b, ONAK12]. RRE [BNS11]. Rule [ARASGAS11, ASARAE12, ONAK12, WWCC07, YCWL01]. Rule-Based [ASARAE12, WWCC07].

S [KL04]. S-clause [KL04]. Salient [CTL00]. Scaffolding [Wan04]. Scale [ZZZL11]. Scenarios [PWK02]. Schema [IA09, LH08]. Scripts [SBN^{+}05]. Search [CLTL02, CC02, CHY03, FI00]. Searching [IA09]. Security [LJAL02]. Segmentation [CC02, HAL09, KL04, LS02a, MBBS05, SS07, WWCC07, WHZ08]. Selected [AK09, KL08, Li09]. Selection [BL02, CLLC08, HHA08, LYK03, LRKL05]. Self [CZZ^{+}05, LC12]. Self-other [LC12]. Self-Partition [CZZ^{+}05]. Semantic [BC06a, CJL12, DC09, GW12, Had07, Hsi11, LH08, YYW01]. Semantics [Lu08, SMB04]. Semi [Yoo02, ZHW08]. Semi-Deterministic [Yoo02]. Semi-Supervised [ZHW08]. Sense [HKAH11, Hsi11, LYK03, WH11, ZMZ12]. Sentence [BL02, LWZF11, SJM05, ZQCL11]. Sentence-based [LWZF11]. Sentences [BNS11, Fai06, ITM^{+}07, NIA06, WWCC07]. Sentiment [TKWL05, XWW08, YLC08, ZHW08, ZQCL11]. Sequences [GW06]. Services [Li00, WL11]. Set [And06, LP96, Lun93]. Shame [LC12]. Shape [PHO02, Wak02]. sharing [PLYZ95]. Short [Kuo11]. SIGHAN [KL08]. Signal [KLC^{+}00]. Signatures [LYY01]. Similarity [ZZ11]. Site [Kwo03]. Sixth [KL08]. Size [HLY^{+}01]. Sketch [GAH08, ŠHHH08]. Smoothing [KKC04]. Soaking [HC12]. Social [SS11]. Solution [LZ03]. Song [XWW08]. Source [GAH08]. Space [CLWK01, XWW08]. Spacing [KKC04]. Speaker [LTB08, Yon06]. Special [AS09a, Ano04a, BC06a, CD03, FZ03, FKL02, KL08, LL01, Lee04, Li08, Li09, Lu08, MW01, SM00, SM02, Won05, iZL07]. Speech [AS09b, BS11, DAZF05, IKK^{+}09, KLC^{+}00, KCL03, LS02b, MBBS05, PTCC05, RC05, STK07, SW00, WCC00]. Spelling [KKC04, PTCC05]. Spoken [Cha01, WC01]. Spreading [AMLJ03]. SSML [JLP04]. Standard [AA04b]. Standards [AA04b]. State [SS07]. State-of-the-art [SS07]. Statistical [ASARAE12, BS11, GW06, NSNN07, OC03, WWCC07, ZLZ11]. Statistics [ZZ11]. Stemming [AS04]. Stenography [TLSC06]. Stenography-To-Text [TLSC06]. Stochastic [KKC04, TM00]. Stock [WXX^{+}08]. Storing [LJAL02]. Strategies [SS01]. Strategy [HLBL00, LKL01]. String [IFK^{+}02, IKK^{+}09]. Stroke [PHO02, SSH05]. Structural [HAL09, KZCJLL03, SNT09]. Structure [CJL12, HASE11, HAH12, Hsi11, SJM05]. Structures [BC06a, DAZF05]. Structuring [RL06]. STT [TLSC06]. Studies [GW06, WXL06]. Study [GAH08, HL12, HKAH11, HC12, LH08, NIA06, WHZ08, WH11, ZZZL11].

Subjectivity [KJLL09]. Sum [HCR03]. Summarization [CTL00, HASE11, LWLW06, SJM05]. Summary [NH05]. Supervised [ZHW08]. Support [AA04a, AKAAS11, HCK^{+}02]. Swapping [CSC06]. Symbols [TCC02]. Symposium [Lee04]. synonymy [GW12]. Syntactic [BC06a, KKSL01, NSNN07, ZQCL11]. Synthesis [WLW03]. Synthetic [LAM08]. System [AA04a, And06, BC06a, BC06b, Cha01, CBA07, CHY03, CHC08, EB09, HH09, KKSL01, Kuo11, LS02b, LWY00, LKL01, LD04, LRKL05, LTT02, LS06, LJAL02, MZA11, MZA12, MBBS05, ML02, TCC02, WLW03, Yon06]. Systems [AKAAS11, BS11, HCK^{+}02, LD04, MW01, YYW01].

Tags [KCL03, SS11]. Taiwan [HL12]. Tamil [SNT09]. Task [BNS11]. Technique [AA04a, HASE11, HC01, LTZ^{+}01]. Techniques [AK09, DSL^{+}02]. Technologies [Sak05, Won05, iZL07]. Technology [Ano04a, FZ03, WLJ03]. Temporal [CAM08, HZLL09, LWY00, LWY03, LWLW06, WLYZ02, WXLY05]. Term [AS04, CLWK01]. Text [AS04, ARASGAS11, BL02, CZZ^{+}05, CAM08, FLW07, HASE11, HAL09, HCR03, LJAL02, MBBS05, RL06, SNT09, SJM05, SCH11, TLSC06, ZZZL11]. Text-To-Speech [MBBS05]. Texts [IA09]. Textual [CTL00]. Thai [AMLJ03, PTCC05, SS01]. their [CPJ93]. Themes [HSBAS12]. Theory [LTB08]. Three [LAM08]. Three-Character [LAM08]. Time [MMFA00, HCK^{+}02]. Tolerant [Ish02]. Tool [KLC^{+}00]. Tools [MZA11, MZA12, Wan04]. Top [AZ04]. Top-Level [AZ04]. Topic [BL02, HASE11, HZLL09, LS02a, MC09]. Topic-Based [MC09]. Topics [iZL07]. Toponym [WL11]. Tracking [LQLY07].

Transfer [KZCJLL03, LHM^{+}00].

Transformation

[NSNN07, VNL02, Wak02, WCC01]. Transformation-based [WCC01]. Transitive [HAH12]. Translation [CYL07, FI00, GW06, KNL05, KNL06, KZCJLL03, KLL04, KY09, LYK03, LRKL05, ML02, NSNN07, Sak01, SRMB04, ZLZ11]. Translational [Tsu02]. Translator [Sah05]. Transliterated [OC03, ŠHHH08]. Transliterating [GW06]. Transliteration [ENB07, JNKL08, KK05]. Transliterations [LKSL08]. Transmission [HC01]. Tree [HCR03]. Trees [LS02b]. Trend [WXX^{+}08]. Truncation [HC02]. TV [Li00]. TV-Like [Li00]. Two [HHA08, HKAH11, KC01, ZZZL11]. Type [CAM08].

Understanding [BL02, LTZ^{+}01]. Unit [LW00]. Units [ML02]. Universities [DSL^{+}02]. University [CAL^{+}00, SM00]. Unknown [LW00, WL05]. Unregistered [WL11]. Unvocalized [IA09]. Upon [TCC02]. Use [MZA11, MZA12]. Used [Yon06]. Using [AK09, AA04b, ARASGAS11, BS11, CC01, CHY03, CZZ^{+}05, CK02, DSL^{+}02, EB08, FK03, GW06, GAH08, HG83, IA09, KASAZS04, KK05, KCL03, Kim06a, KJLL09, KY09, Kuo11, KKC04, LW00, LS02b, LJL11, LHM^{+}00, LKSL08, LS06, MC09, MH00, PTCC05, Sak01, Sar09, SNT09, SW00, TM00, TT00, Wak02, WCC00, WWCC07, YLC08, Yoo02, ZMI05, ZHW08, LP96].

Vector [CLWK01, VNL02, XWW08]. Vector-to-Image [VNL02]. Vectorized [CHC08]. Verb [HAH12, IIK^{+}11]. Verbs [BM08, GW12, HHA08, HC12, LH08]. Verification [CK02, LYY01]. via [LJAL02]. Virtual [DSL^{+}02, Jin00, SM00, WLJ03, YYW01, CCLC04]. Vocabulary [DAZF05, LS06, WHZ08]. Vocalized [IA09].

Vocalized/Unvocalized [IA09]. Volume [Ano03, Ano04b, Ano05, Ano06, Ano07, Ano08, Ano09, Ano11]. Vowel [AMLJ03, LTB08]. Vowels [AMLJ03]. VPL [CCLC04]. vs [KNL06].

WAP [SM02]. Wavelet [SW00]. Ways [HL12]. Web [Ano04a, CHY03, FZ03, KOK03, Kwo03, LTZ^{+}01, PWK02, SL03, SM02, WCC00, WLJ03, WL11, ZMWP04]. Web-Based [WLJ03, ZMWP04, Ano04a, FZ03, SL03]. Weblog [YLC08]. Weighted [CLWK01, HCR03]. Weighting [SW00]. West [AS09b]. Width [CC01]. Window [HLY^{+}01]. Word [DC09, GAH08, JNKL08, KC01, KY09, KKC04, LW00, LYK03, LS06, ML02, ŠHHH08, TM00, Tsu02, WHSN04, WL05, WHZ08, ZZ11, ZMZ12]. Word-alignment [KY09]. Word-Spacing [KKC04]. Wordlist [Kwo01]. Wordnet [Hsi11]. Words [HY07, KC01, KNL06, LC12, LAM08, MH00, OC03]. Workflow [WLR01]. Working [HCK^{+}02]. Workshop [KL08, Lu08]. Writing [BC06a]. Written [NH05]. WSD [QLZ^{+}08]. WWW [KSH00].

XML [PWK02]. XML-based [PWK02].

References

Abuleil:2004:NTS

[AA04a] Saleem Abuleil and Khalid Alsamara. New technique to support Arabic noun morphology: Arabic Noun Classifier System (ANCS). International Journal of Computer Processing of Oriental Languages (IJCPOL), 17(2): 97–??, June 2004. CODEN ???? ISSN 0219-4279.

Al-Anzi:2004:PAS

[AA04b]  Fawaz S. Al-Anzi.  Pro-

ducing algorithmically standard romanization of Arabic names using hints from non-standards. International Journal of Computer Processing of Oriental Languages (IJCPOL), 17(3):165– ??, September 2004. CODEN ???? ISSN 0219-4279.

Abandah:2009:AHA

[AK09] Gheith A. Abandah and Mohammed Z. Khedher. Analysis of handwritten Arabic letters using selected feature extraction techniques. International Journal of Computer Processing Of Languages (IJCPOL), 22(1):49–73, March 2009. CO-DEN ???? ISSN 2151-3848.

Abuleil:2011:TBE

[AKAAS11] Saleem Abuleil, Khalid Khanfar, Khalid Alsamara, and Hani Abu-Salem. Toward building events database to support information retrieval systems. International Journal of Computer Processing Of Languages (IJCPOL), 23 (4):409–??, December 2011. CODEN ???? ISSN 2151-3848.

Ahkuputra:2003:AMV

[AMLJ03] Visarut Ahkuputra, Ekkarit Maneenoi, Sudaporn Luksaneeyanawin, and Somchai Jitapunkul. Acoustic modelling of vowel articulation on the nine Thai spreading vowels.

International Journal of Computer Processing of Oriental Languages (IJCPOL), 16(3): 171–??, September 2003. CO-DEN ???? ISSN 0219-4279.

Andres:2006:DAI

[And06] Paul Andres. Developing an appropriate icon set for a Mandarin Chinese augmentative communication system. International Journal of Computer Processing of Oriental Languages (IJCPOL), 19(4): 275–283, December 2006. CO-DEN ???? ISSN 0219-4279.

Anonymous:2000:AI

[Ano00] Anonymous. Author index.

International Journal of Computer Processing of Oriental Languages (IJCPOL), 13(4): 377–??, 2000. CODEN ???? ISSN 0219-4279.

Anonymous:2001:AI

[Ano01] Anonymous. Author index.

International Journal of Computer Processing of Oriental Languages (IJCPOL), 14(4): 375–??, 2001. CODEN ???? ISSN 0219-4279.

Anonymous:2002:AI

[Ano02] Anonymous. Author index.

International Journal of Computer Processing of Oriental Languages (IJCPOL), 15(4): 517–??, 2002. CODEN ???? ISSN 0219-4279.

Anonymous:2003:AIV

[Ano03] Anonymous. Author index volume 16 (2003). International Journal of Computer Processing of Oriental Languages (IJCPOL), 16(4):311– ??, December 2003. CODEN ???? ISSN 0219-4279.

Anonymous:2004:GEI

[Ano04a] Anonymous. Guest Editors’ introduction: Special issue: Web-based educational technology in the Asia–Pacific Rim (Part II). International Journal of Computer Processing of Oriental Languages (IJCPOL), 17(1):1–??, March 2004. CODEN ???? ISSN 0219-4279.

Anonymous:2004:VAI

[Ano04b]  Anonymous.  Volume 17

(2004) author index. International Journal of Computer Processing of Oriental Languages (IJCPOL), 17(4):299– ??, December 2004. CODEN ???? ISSN 0219-4279.

Anonymous:2005:AIV

[Ano05] Anonymous. Author index volume 18 (2005). International Journal of Computer Processing of Oriental Languages (IJCPOL), 18(4):303– ??, December 2005. CODEN ???? ISSN 0219-4279.

Anonymous:2006:AIV

[Ano06] Anonymous. Author index, volume 19 (2006). International Journal of Computer Processing of Oriental Languages (IJCPOL), 19(4):301– 302, December 2006. CODEN ???? ISSN 0219-4279.

Anonymous:2007:AIV

[Ano07] Anonymous. Author index, volume 20 (2007). International Journal of Computer Processing of Oriental Languages (IJCPOL), 20(4):311– 312, December 2007. CODEN ???? ISSN 0219-4279.

Anonymous:2008:AIV

[Ano08] Anonymous. Author index, volume 21 (2008). International Journal of Computer Processing Of Languages (IJCPOL), 21(4):363– 364, December 2008. CODEN ???? ISSN 2151-3848.

Anonymous:2009:AIV

[Ano09] Anonymous. Author index, volume 22 (2009). International Journal of Computer Processing Of Languages (IJCPOL), 22(4):381– 382, December 2009. CODEN ???? ISSN 2151-3848.

Anonymous:2011:AIV

[Ano11] Anonymous. Author index, volume 23 (2011). International Journal of Computer Processing Of Languages (IJCPOL), 23(4):441– ??, December 2011. CODEN ???? ISSN 2151-3848.

Al-Radaideh:2011:AAT

[ARASGAS11] Qasem A. Al-Radaideh, Emad M.

Al-Shawakfa, Abdullah S. Ghareb, and Hani Abu-Salem. An approach for Arabic text categorization using association rule mining. International Journal of Computer Processing Of Languages (IJCPOL), 23(1):81– 106, March 2011. CO-DEN ???? ISSN 2151-

3848.  URL http://www.
worldscinet.com/ijcpol/
23/preserved-docs/2301/
S179384061100222X.pdf.

Abu-Salem:2004:CSG

[AS04] Hani Abu-Salem. Comparison of stemming and N-gram matching for term conflation in Arabic text. International Journal of Computer Processing of Oriental Languages (IJCPOL), 17(2):61– ??, June 2004. CODEN ???? ISSN 0219-4279.

Abu-Salem:2009:GEI

[AS09a] Hani Abu-Salem. Guest Editor’s introduction: Special issue: Arabic natural language processing. International Journal of Computer Processing Of Languages (IJCPOL), 22(4):257– 267, December 2009. CODEN ???? ISSN 2151-3848.

Alotaibi:2009:EMW

[AS09b] Yousef Ajami Alotaibi and Sid-Ahmed Selouani. Evaluating the MSA West Point Speech Corpus. International Journal of Computer Processing Of Languages (IJCPOL), 22(4):285– 304, December 2009. CODEN ???? ISSN 2151-3848.

Al-Shawakfa:2012:ERB

[ASARAE12] Emad M. Al-Shawakfa, Qasem A.

Al-Radaideh, and Ahmed F. Al-Eroud. Evaluating rulebased and statistical filters to detecting Arabic e-mail alert messages. International Journal of Computer Processing Of Languages (IJCPOL), 24 (2):111–??, June 2012. CO-DEN ???? ISSN 2151-3848.

Abu-Salem:2006:EAC

[ASC06] Hani Abu-Salem and Philip Ka-Fai Chan. English– Arabic cross-language information retrieval based on parallel documents. International Journal of Computer Processing of Oriental Languages (IJCPOL), 19(1):21– ??, March 2006. CODEN ???? ISSN 0219-4279. URL http:

//ejournals.wspc.com.sg/
ijcpol/19/1901/S0219427906001372.
html.

Al-Taani:2011:APA

[ATAAAS11] Ahmad T. Al-Taani, Noor Aldeen K. Al-Awad, and Hani Abu-Salem. An adaptive parser for Arabic language processing. International Journal of Computer Processing Of Languages (IJCPOL), 23(1):67–80, March 2011. CO-DEN ???? ISSN 2151-

3848.  URL http://www.
worldscinet.com/ijcpol/
23/preserved-docs/2301/
S1793840611002218.pdf.

Al-Taani:2009:ANC

[ATWD09] Ahmad T. Al-Taani, Sana A. Wedian, and Omar M. Darwish. Arabic numerals checker: Checking agreement between numerals and counted objects in the Arabic language. International Journal of Computer Processing Of Languages (IJCPOL), 22 (4):341–357, December 2009. CODEN ???? ISSN 2151-3848.

Al-Zoman:2004:TLA

[AZ04] Abdulaziz H. Al-Zoman. Toplevel Arabic domain names. International Journal of Computer Processing of Oriental Languages (IJCPOL), 17(2): 83–??, June 2004. CODEN ???? ISSN 0219-4279.

Baker:2006:GEI

[BC06a] Bruce R. Baker and Shi-Kuo Chang. Guest Editors’ introduction: Special issue: Syntactic and semantic structures of the Mandarin Chinese language and writing system for augmentative communication (AC). International Journal of Computer Processing of Oriental Languages (IJCPOL), 19(4):223– 224, December 2006. CODEN ???? ISSN 0219-4279.

Baker:2006:MLS

[BC06b] Bruce R. Baker and Shi-Kuo Chang. A Mandarin language system in augmentative and alternative communication (AAC). International Journal of Computer Processing of Oriental Languages (IJCPOL), 19(4):225– 237, December 2006. CODEN ???? ISSN 0219-4279.

Bae:2002:MDT

[BL02] Jae-Hak J. Bae and Jong-Hyeok Lee. Mid-depth text understanding by abductive chains for topic sentence selection. International Journal of Computer Processing of Oriental Languages (IJCPOL), 15(3):341– ??, September 2002. CODEN ???? ISSN 0219-4279.

Belkredim:2008:DDA

[BM08] Fatma Zohra Belkredim and Farid Meziane. DEAR-ONTO: A DErivational ARabic ontology based on verbs. International Journal of Computer Processing Of Languages (IJCPOL), 21(3):279– 291, September 2008. CO-DEN ???? ISSN 2151-3848.

Bach:2011:RTT

[BNS11] Ngo Xuan Bach, Le Minh Nguyen, and Akira Shimazu. RRE task: The task of recognition of requisite part and effectuation part in law sentences. International Journal of Computer Processing Of Languages (IJCPOL), 23 (2):109–130, June 2011. CO-DEN ???? ISSN 2151-3848.

Bahrani:2011:BSL

[BS11] Mohammad Bahrani and Hossein Sameti. Building statistical language models for Persian continuous speech recognition systems using the Peykare Corpus. International Journal of Computer Processing Of Languages (IJCPOL), 23(1):1–20, March 2011. CO-DEN ???? ISSN 2151-

                                        3848.  URL http://www.
                                        worldscinet.com/ijcpol/
                                        23/preserved-docs/2301/
                                        S1793840611002188.pdf.

Chang:2000:MU

[CAL^{+}00] S. K. Chang, T. Arndt, S. Levialdi, A. C. Liu, J. Ma, T. Shih, and G. Tortora. Macro university. International Journal of Computer Processing of Oriental Languages (IJCPOL), 13(3):205– ??, 2000. CODEN ???? ISSN 0219-4279.

Cheng:2008:EET

[CAM08] Yuchang Cheng, Masayuki Asahara, and Yuji Matsumoto. Effect of event type classification for temporal relation identification in Chinese text. International Journal of Computer Processing Of Languages (IJCPOL), 21 (2):151–167, June 2008. CO-DEN ???? ISSN 2151-3848.

Charfi:2007:ISD

[CBA07] Moncef Charfi, Wafa Boussellaa, and Mohamed Adel Alimi. An intelligent system for digitization and processing of historical Arabic documents. International Journal of Computer Processing of Oriental Languages (IJCPOL), 20(4): 259–288, December 2007. CO-DEN ???? ISSN 0219-4279.

Chang:2001:BLL

[CC01] Chin-Chen Chang and Guu-In Chen. Bi-level lossless calligraphy compression using location propagation and width inheritance. International Journal of Computer Processing of Oriental Languages (IJCPOL), 14(4):309– ??, 2001. CODEN ???? ISSN 0219-4279.

Chen:2002:NSE

[CC02] Tung-Shou Chen and Ming-Chiang Chen. A new search engine for Chinese document image retrieval based on Chinese character segmentation features. International Journal of Computer Processing of Oriental Languages (IJCPOL), 15(4):417– ??, 2002. CODEN ???? ISSN 0219-4279.

Chan:2004:VOD

[CCLC04] Alvin T. S. Chan, Jiannong Cao, Chi-Kin Liu, and Weidong Cao. VPL: An online distance learning platform for Virtual Programming Laboratory. International Journal of Computer Processing of Oriental Languages (IJCPOL), 17(1):41–??, March 2004. CO-DEN ???? ISSN 0219-4279.

Chao:2003:GEI

[CD03] Ming Chao and Chyi-Ren Dow. Guest Editors’ introduction: Special issue: Best papers from ICCLC 2002. International Journal of Computer Processing of Oriental Languages (IJCPOL), 16(2): 103–??, June 2003. CODEN ???? ISSN 0219-4279.

Chang:2001:TOB

[Cha01]  Chao-Huang Chang.  To-

wards ontology-based multilingual spoken dialogue system. International Journal of Computer Processing of Oriental Languages (IJCPOL), 14(1):1–15, 2001. CODEN ???? ISSN 0219-4279.

Chiang:2008:VCF

[CHC08] John Y. Chiang, Chung-Hsun Hsieh, and Shuenn-Ren Cheng. A vectorized Chinese font personalizing system. International Journal of Computer Processing Of Languages (IJCPOL), 21(3):259– 278, September 2008. CO-DEN ???? ISSN 2151-3848.

Chen:2003:UDM

[CHY03] Yu-Ru Chen, Ming-Chuan Hung, and Don-Lin Yang. Using data mining to construct an intelligent Web search system. International Journal of Computer Processing of Oriental Languages (IJCPOL), 16(2):143–??, June 2003. CO-DEN ???? ISSN 0219-4279.

Chen:2012:BCS

[CJL12] Bo Chen, Donghong Ji, and Chen Lv. Building a Chinese semantic resource based on feature structure. International Journal of Computer Processing Of Languages (IJCPOL), 24(1):95– ??, March 2012. CODEN ???? ISSN 2151-3848.

Cho:2002:VGU

[CK02] Sung-Jung Cho and Jin H.

Kim. Verification of graphemes using neural networks for HMM-based on-line handwritten Hangul recognition. International Journal of Computer Processing of Oriental Languages (IJCPOL), 15(2): 133–??, June 2002. CODEN ???? ISSN 0219-4279.

Chen:2002:RP

[CL02] Jason Jen-Yen Chen and Chun-Yi Lin. Regular papers. International Journal of Computer Processing of Oriental Languages (IJCPOL), 15(4): 457–??, 2002. CODEN ???? ISSN 0219-4279.

Cui:2008:ASC

[CLLC08] Gaoying Cui, Qin Lu, Wenjie Li, and Yirong Chen. Attributes selection in Chinese ontology acquisition with FCA. International Journal of Computer Processing Of Languages (IJCPOL), 21(1):77– 95, March 2008. CODEN ???? ISSN 2151-3848.

Chen:2002:DCS

[CLTL02] Tung-Shou Chen, Kuang-Fu Li, Yu-Cheng Tsai, and Wen-Ping Lin. Design of a Chinese search engine based on homophonic Chinese character encoding. International Journal of Computer Processing of Oriental Languages (IJCPOL), 15(4):407– ??, 2002. CODEN ???? ISSN 0219-4279.

Chow:2001:HTI

[CLWK01] Ken C. W. Chow, Robert W. P. Luk, K. F. Wong, and K. L. Kwok. Hybrid term indexing for weighted Boolean and vector space models. International Journal of Computer Processing of Oriental Languages (IJCPOL), 14(2): 133–151, 2001. CODEN ???? ISSN 0219-4279.

Chan:1993:NNT

[CPJ93] L. S.-C. Chan, Hean-Lee Poh, and Teo Jasic. Neural networks and their applications. Computer Processing of Chinese and Oriental Languages: an international journal of the Chinese Language Computer Society, 7(2):133–66, December 1993. CODEN CPCLE6. ISSN 0715-9048.

Choi:2006:IBP

[CSC06] Yong-Seok Choi, Ji-Ae Shin, and Key-Sun Choi. Identification of boundaries in parallel noun phrases: A probabilistic swapping model. International Journal of Computer Processing of Oriental Languages (IJCPOL), 19(2– 3):109–132, June/September 2006. CODEN ???? ISSN 0219-4279.

Chan:2000:ISI

[CTL00] Samuel W. K. Chan, Benjamin K. T’sou, and Tom B. Y. Lai. An investigation of salient information in textual summarization. International Journal of Computer Processing of Oriental Languages (IJCPOL), 13(4):265– ??, 2000. CODEN ???? ISSN 0219-4279.

Chen:2007:MEM

[CYL07] Yin Chen, Muyun Yang, and Sheng Li. Maximum entropy model for examplebased machine translation. International Journal of Computer Processing of Oriental Languages (IJCPOL), 20(2– 3):101–113, June/September 2007. CODEN ???? ISSN 0219-4279.

Chen:2005:IDD

[CZZ^{+}05] Wenliang Chen, Jingbo Zhu, Muhua Zhu, Li Zhang, and Tianshun Yao. Improving domain dictionary-based text categorization using selfpartition model. International Journal of Computer Processing of Oriental Languages (IJCPOL), 18(3):197– ??, September 2005. CODEN ???? ISSN 0219-4279.

Datta:2005:EMS

[DAZF05] S. Datta, M. Al-Zabibi, and O. Farooq. Exploitation of morphological structures in large vocabulary Arabic speech recognition. International Journal of Computer Processing of Oriental Languages (IJCPOL), 18(4):291– ??, December 2005. CODEN ???? ISSN 0219-4279.

Ding:2009:WBC

[DC09] Weiwei Ding and Baobao Chang. Word based Chinese semantic role labeling with semantic chunking. International Journal of Computer Processing Of Languages (IJCPOL), 22(2– 3):133–154, June/September 2009. CODEN ???? ISSN 2151-3848.

Dow:2002:VLM

[DSL^{+}02] C. R. Dow, C. C. Shen, J. H. Lin, S. C. Chen, and C. Y. Lin. A virtual laboratory for macro universities using mobile agent techniques. International Journal of Computer Processing of Oriental Languages (IJCPOL), 15(1):1–??, 2002. CODEN ???? ISSN 0219-4279.

Ekbal:2008:NER

[EB08] Asif Ekbal and Sivaji Bandyopadhyay. Named entity recognition in Indian languages using maximum entropy approach. International Journal of Computer Processing Of Languages (IJCPOL), 21(3):205–237, September 2008. CODEN ???? ISSN 2151-3848.

Ekbal:2009:MNS

[EB09] Asif Ekbal and Sivaji Bandyopadhyay. A multiengine NER system with context pattern learning and post-processing improves system performance. International Journal of Computer Processing Of Languages (IJCPOL), 22(2–3): 171–204, June/September 2009. CODEN ???? ISSN 2151-3848.

Ekbal:2007:NET

[ENB07] Asif Ekbal, Sudip Kumar Naskar, and Sivaji Bandyopadhyay. Named entity transliteration. International Journal of Computer Processing of Oriental Languages (IJCPOL), 20(4):289– 310, December 2007. CODEN ???? ISSN 0219-4279.

Faiz:2006:IRS

[Fai06] Rim Faiz. Identifying relevant sentences in news articles for event information extraction. International Journal of Computer Processing of Oriental Languages (IJCPOL), 19(1):1–??, March 2006. CODEN ???? ISSN 0219-4279. URL http:

//ejournals.wspc.com.sg/
ijcpol/19/1901/S0219427906001384.
html.

Fujii:2000:CLI

[FI00] A. Fujii and T. Ishikawa. Cross-language information retrieval based on query keyword translation: An Internet search application. International Journal of Computer Processing of Oriental Languages (IJCPOL), 13(1):1–??, 2000. CODEN ???? ISSN 0219-4279.

Fong:2003:ELU

[FK03] Joseph Fong and Irene Kwan. Effective e-learning by using HCI and interactivity on data modeling. International Journal of Computer Processing of Oriental Languages (IJCPOL), 16(4):293–??, December 2003. CODEN ???? ISSN 0219-4279.

Fujisawa:2002:GEI

[FKL02] Hiromichi Fujisawa, Jin H. Kim, and Cheng Lin Liu. Guest Editors’ introduction: Special issue on oriental character recognition. International Journal of Computer Processing of Oriental Languages (IJCPOL), 15(2):115– ??, June 2002. CODEN ???? ISSN 0219-4279.

Fu:2007:AEA

[FLW07] Guohong Fu, Kang-Kwong Luke, and Jonathan J. Webster. Automatic expansion of abbreviations in Chinese news text: A hybrid approach. International Journal of Computer Processing of Oriental Languages (IJCPOL), 20(2– 3):165–179, June/September 2007. CODEN ???? ISSN 0219-4279.

Fong:2003:GEI

[FZ03] Joseph Fong and Wanlei Zhou. Guest Editors’ introduction: Special issue: Webbased educational technology in the Asia–Pacific Rim (Part I). International Journal of Computer Processing of Oriental Languages (IJCPOL), 16(4):241–??, December 2003. CODEN ???? ISSN 0219-4279.

Gong:2008:CWS

[GAH08] Shu-Ping Gong, Kathleen Ahrens, and Chu-Ren Huang. Chinese word sketch and mapping principles: A corpusbased study of conceptual metaphors using the BUILD-ING source domain. International Journal of Computer Processing Of Languages (IJCPOL), 21(1):3–17, March 2008. CODEN ???? ISSN 2151-3848.

Gao:2006:ESU

[GW06] Wei Gao and Kam-Fai Wong.

Experimental studies using statistical algorithms on transliterating phoneme sequences for English–Chinese name translation. International Journal of Computer Processing of Oriental Languages (IJCPOL), 19(1):63– ??, March 2006. CODEN ???? ISSN 0219-4279. URL http:

                                        //ejournals.wspc.com.sg/
                                        ijcpol/19/1901/S0219427906001396.
                                        html.

Gong:2012:CSP

[GW12] Shu-Ping Gong and Pei-Yan Wu. Collocation, semantic prosody, and near-synonymy: The HELP verbs in Mandarin Chinese. International Journal of Computer Processing Of Languages (IJCPOL), 24 (1):3–??, March 2012. CO-DEN ???? ISSN 2151-3848.

Haddad:2007:SRA

[Had07] Bassam Haddad. Semantic representation of Arabic: A logical approach towards compositionality and generalized Arabic quantifiers. International Journal of Computer Processing of Oriental Languages (IJCPOL), 20(1):37– 52, March 2007. CODEN ???? ISSN 0219-4279.

Hong:2012:EST

[HAH12] Jia-Fei Hong, Kathleen Ahrens, and Chu-Ren Huang. Event structure of transitive verb: A MARVS perspective. International Journal of Computer Processing Of Languages (IJCPOL), 24(1):37– ??, March 2012. CODEN ???? ISSN 2151-3848.

Hanmandlu:2009:SHH

[HAL09] M. Hanmandlu, Pooja Agrawal, and Brejesh Lall. Segmentation of handwritten Hindi text: A structural approach.

International Journal of Computer Processing Of Languages (IJCPOL), 22(1):1–20, March 2009. CODEN ???? ISSN 2151-3848.

Hammo:2011:HAT

[HASE11] Bassam H. Hammo, Hani Abu-Salem, and Martha W. Evens. A hybrid Arabic text summarization technique based on text structure and topic identification. International Journal of Computer Processing Of Languages (IJCPOL), 23(1): 39–65, March 2011. CO-DEN ???? ISSN 2151-

3848.  URL http://www.
worldscinet.com/ijcpol/
23/preserved-docs/2301/
S1793840611002206.pdf.

Hung:2001:NPI

[HC01] Kuo-Lung Hung and Chin-Chen Chang. A novel progressive image transmission technique for Chinese calligraphy. International Journal of Computer Processing of Oriental Languages (IJCPOL), 14(4): 293–??, 2001. CODEN ???? ISSN 0219-4279.

Hung:2002:EPC

[HC02] Kuo-Lung Hung and Chin-Chen Chang. Error prevention and concealment for Chinese calligraphy based on block truncation coding. International Journal of Computer Processing of Oriental Languages (IJCPOL), 15(1):

101–??, 2002. CODEN ???? ISSN 0219-4279.

Hsu:2012:CBS

[HC12] You-Fen Hsu and Siaw-Fong Chung. A corpus-based study of Mandarin soaking verbs pào and jìn. International Journal of Computer Processing Of Languages (IJCPOL), 24 (1):17–??, March 2012. CO-DEN ???? ISSN 2151-3848.

He:2002:WMR

[HCK^{+}02] Aiguo He, Zixue Cheng, Akio Koyama, Tong Jun Huang, Jerold A. Dehart, and Yue Zhao. A working model for real-time interactive distance education support systems. International Journal of Computer Processing of Oriental Languages (IJCPOL), 15(1): 19–??, 2002. CODEN ???? ISSN 0219-4279.

Hwang:2003:WPS

[HCR03] Young-Sook Hwang, Hoojung Chung, and Hae-Chang Rim. Weighted probabilistic sum model based on decision tree decomposition for text chunking. International Journal of Computer Processing of Oriental Languages (IJCPOL), 16(1):1–??, March 2003. CO-DEN ???? ISSN 0219-4279.

Hobby:1983:UMD

[HG83] John D. Hobby and G. Gu. Using Metafont to design Chinese characters. Computer Processing of Chinese and Oriental Languages: an international journal of the Chinese Language Computer Society, 1(1):4–23, July 1983. CODEN CPCLE6. ISSN 0715-9048.

Haddar:2009:ERS

[HH09] Kais Haddar and Abdelmajid Ben Hamadou. An ellipsis resolution system for the Arabic language. International Journal of Computer Processing Of Languages (IJCPOL), 22(4):359– 380, December 2009. CODEN ???? ISSN 2151-3848.

Hsieh:2012:GEI

[HH12] Shu-Kai Hsieh and Jia-Fei Hong. Guest Editors’ introduction. International Journal of Computer Processing Of Languages (IJCPOL), 24 (1):1–??, March 2012. CO-DEN ???? ISSN 2151-3848.

Hong:2008:ESC

[HHA08] Jia-Fei Hong, Chu-Ren Huang, and Kathleen Ahrens. Event selection and coercion of two verbs of ingestion: A MARVS perspective. International Journal of Computer Processing Of Languages (IJCPOL), 21(1):31–42, March 2008. CO-DEN ???? ISSN 2151-3848.

Hill:2006:AAC

[Hil06] Katya Hill. Augmentative and alternative communication (AAC) research and development: The challenge of evidence-based practice. International Journal of Computer Processing of Oriental Languages (IJCPOL), 19(4): 249–262, December 2006. CO-DEN ???? ISSN 0219-4279.

Hong:2011:SPS

[HKAH11] Jia-Fei Hong, Sue-Jin Ker, Kathleen Ahrens, and Chu-Ren Huang. Sense prediction study: Two corpus-driven linguistic approaches. International Journal of Computer Processing Of Languages (IJCPOL), 23(3):229– ??, September 2011. CODEN ???? ISSN 2151-3848.

Her:2012:CMW

[HL12] One-Soon Her and Wan-Jun Lai. Classifiers: The many ways to profile ‘one’ — a case study of Taiwan Mandarin. International Journal of Computer Processing Of Languages (IJCPOL), 24 (1):79–??, March 2012. CO-DEN ???? ISSN 2151-3848.

Hwang:2000:RSM

[HLBL00] Y.-G. Hwang, H.-Y. Lee, W.- J. Bae, and Y.-S. Lee. Resolution strategy of morphological ambiguity for Korean parsing. International Journal of Computer Processing of Oriental Languages (IJCPOL), 13(2): 97–??, 2000. CODEN ???? ISSN 0219-4279.

Hung:2001:DCW

[HLY^{+}01] Kei Yuen Hung, Robert Luk, Daniel Yeung, Korris Chung, and Wenhuo Shu. Determination of context window size. International Journal of Computer Processing of Oriental Languages (IJCPOL), 14(1): 71–80, 2001. CODEN ???? ISSN 0219-4279.

Hammo:2012:CAI

[HSBAS12] Bassam H. Hammo, Azzam Sleit, Aladdin Baarah, and Hani Abu-Salem. A computational approach for identifying Quranic themes. International Journal of Computer Processing Of Languages (IJCPOL), 24(2):189–??, June 2012. CO-DEN ???? ISSN 2151-3848.

Hsieh:2011:SSC

[Hsi11] Shu-Kai Hsieh. Sense structure in cube: Lexical semantic representation in Chinese wordnet. International Journal of Computer Processing Of Languages (IJCPOL), 23 (3):243–??, September 2011. CODEN ???? ISSN 2151-3848.

Hsu:1992:CCP

[HW92] Loke Soo Hsu and Zhibiao Wu. Chinese character prediction by recurrent network. Computer Processing of Chinese and Oriental Languages: an international journal of the Chinese Language Computer Society, 6(2):179–194, December 1992. CODEN CPCLE6. ISSN 0715-9048.

Haddad:2007:DCN

[HY07] Bassam Haddad and Mustafa Yaseen. Detection and correction of non-words in Arabic: A hybrid approach. International Journal of Computer Processing of Oriental Languages (IJCPOL), 20(4): 237–257, December 2007. CO-DEN ???? ISSN 0219-4279.

Hong:2009:TTM

[HZLL09] Yu Hong, Yu Zhang, Ting Liu, and Sheng Li. A temporal topic model for new event detection. International Journal of Computer Processing Of Languages (IJCPOL), 22 (1):99–130, March 2009. CO-DEN ???? ISSN 2151-3848.

Ibrahim:2009:SVU

[IA09] Farid Ibrahim and Jalal Omer Atoum. Searching vocalized/unvocalized Arabic texts using an improved coding schema. International Journal of Computer Processing Of Languages (IJCPOL), 22 (4):269–283, December 2009. CODEN ???? ISSN 2151-3848.

Ikeda:2002:CFG

[IFK^{+}02] Hisashi Ikeda, Naohiro Furukawa, Masashi Koga, Hiroshi Sako, and Hiromichi Fujisawa. A context-free grammar-based language model for string recognition. International Journal of Computer Processing of Oriental Languages (IJCPOL), 15(2):149– ??, June 2002. CODEN ???? ISSN 0219-4279.

Izumi:2011:PJL

[IIK^{+}11] Tomoko Izumi, Kenji Imamura, Genichiro Kikui, Atsushi Fujita, and Satoshi Sato. Paraphrasing Japanese light verb constructions: Towards the normalization of complex predicates. International Journal of Computer Processing Of Languages (IJCPOL), 23(2):147–167, June 2011. CODEN ???? ISSN 2151-3848.

Ito:2009:DJS

[IKK^{+}09] Akinori Ito, Hiroaki Kinno, Masaharu Katoh, Tetsuo Kosaka, and Masaki Kohda. Dictation of Japanese speech based on Kana and Kanji character string. International Journal of Computer Processing Of Languages (IJCPOL), 22(1):75–98, March 2009. CO-DEN ???? ISSN 2151-3848.

Ishitani:2002:MBI

[Ish02]  Yasuto Ishitani.  Model-

based information extraction method tolerant of OCR errors for document images. International Journal of Computer Processing of Oriental Languages (IJCPOL), 15(2): 165–??, June 2002. CODEN ???? ISSN 0219-4279.

Ikehara:2007:PDD

[ITM^{+}07] Satoru Ikehara, Masato Tokuhisa, Jin’Ichi Murakami, Masashi Saraki, Masahiro Miyazaki, and Naoshi Ikeda. Pattern dictionary development based on non-compositional language model for Japanese compound and complex sentences. International Journal of Computer Processing of Oriental Languages (IJCPOL), 20(2–3):151–163, June/September 2007. CO-DEN ???? ISSN 0219-4279.

Zhang:2007:GEI

[iZL07] in Zhang and Haizhou Li. Guest Editors’ introduction: Special issue: Advanced topics in language technologies. International Journal of Computer Processing of Oriental Languages (IJCPOL), 20 (2–3):75–77, June/September 2007. CODEN ???? ISSN 0219-4279.

Jin:2000:MVE

[Jin00] Qun Jin. A multilingual virtual environment for everycitizen learning communities. International Journal of Computer Processing of Oriental Languages (IJCPOL), 13(3): 223–??, 2000. CODEN ???? ISSN 0219-4279.

Jin:2004:CSD

[JLP04] Lee Hwa Jin, Ho-Joon Lee, and Jong C. Park. Constructing SSML documents with automatically generated intonation information in a combinatory categorial grammar framework. International Journal of Computer Processing of Oriental Languages (IJCPOL), 17(4):223– ??, December 2004. CODEN ???? ISSN 0219-4279.

Jin:2008:ECT

[JNKL08] Chengguo Jin, Seung-Hoon Na, Dong-Il Kim, and Jong-Hyeok Lee. English–Chinese transliteration word pair extraction from parallel corpora. International Journal of Computer Processing Of Languages (IJCPOL), 21(2):169– 182, June 2008. CODEN ???? ISSN 2151-3848.

Jean:2000:PBC

[JT00] En-Yih Jean and Cheng-Huang Tung. A phonemebased Chinese input method with low conflict code rate. International Journal of Computer Processing of Oriental Languages (IJCPOL), 13(4): 333–??, 2000. CODEN ???? ISSN 0219-4279.

Kanaan:2004:CBA

[KASAZS04] Ghassan Kanaan, Riyad Al-Shalabi, Mohammed G. Al-Zamil, and Ahmad A. Saifan. Comparison between ad-hoc retrieval and filtering retrieval using Arabic documents. International Journal of Computer Processing of Oriental Languages (IJCPOL), 17(3): 181–??, September 2004. CO-DEN ???? ISSN 0219-4279.

Kang:2001:TAR

[KC01] Byung-Ju Kang and Key-Sun Choi. Two approaches for the resolution of word mismatch problem caused by English words and foreign words in Korean information retrieval. International Journal of Computer Processing of Oriental Languages (IJCPOL), 14(2): 109–131, 2001. CODEN ???? ISSN 0219-4279.

Kim:2002:RP

[KCL02] Dongseok Kim, Jeongwon Cha, and Gary Geunbae Lee. Regular papers. International Journal of Computer Processing of Oriental Languages (IJCPOL), 15(1):63–??, 2002. CODEN ???? ISSN 0219-4279.

Kim:2003:AGC

[KCL03] Seongyong Kim, Key-Sun Choi, and Kong Joo Lee. Automatic generation of composite labels using part-of-speech tags for parsing Korean. International Journal of Computer Processing of Oriental Languages (IJCPOL), 16(3): 197–??, September 2003. CO-DEN ???? ISSN 0219-4279.

Kim:2006:CRC

[Kim06a] Mi-Young Kim. Classification of relative clauses using easily obtainable features. International Journal of Computer Processing of Oriental Languages (IJCPOL), 19(2– 3):189–199, June/September 2006. CODEN ???? ISSN 0219-4279.

Kim:2006:HMD

[Kim06b] Mi-Young Kim. A hybrid method to determine interclause dependency. International Journal of Computer Processing of Oriental Languages (IJCPOL), 19(2–3): 89–97, June/September 2006. CODEN ???? ISSN 0219-4279.

Kim:2009:CSL

[KJLL09] Jungi Kim, Hun-Young Jung, Yeha Lee, and Jong-Hyeok Lee. Conveying subjectivity of a lexicon of one language into another using a bilingual dictionary and a link analysis algorithm. International Journal of Computer Processing Of Languages (IJCPOL), 22(2– 3):205–218, June/September 2009. CODEN ???? ISSN 2151-3848.

Kang:2005:TUN

[KK05] In-Ho Kang and Gil Chang Kim. Transliteration using a network of phoneme chunks. International Journal of Computer Processing of Oriental Languages (IJCPOL), 18(1): 1–??, March 2005. CODEN ???? ISSN 0219-4279.

Kwon:2004:SKW

[KKC04] Hyuk-Chul Kwon, Mi-Young Kang, and Sung-Ja Choi. Stochastic Korean wordspacing with smoothing using Korean spelling checker. International Journal of Computer Processing of Oriental Languages (IJCPOL), 17(4): 239–??, December 2004. CO-DEN ???? ISSN 0219-4279.

Kim:2001:FRQ

[KKSL01] Harksoo Kim, Kyungsun Kim, Jungyun Seo, and Gary Geunbae Lee. A fast and reliable question-answering system based on predictive answer indexing and lexicosyntactic pattern matching. International Journal of Computer Processing of Oriental Languages (IJCPOL), 14(4): 341–??, 2001. CODEN ???? ISSN 0219-4279.

Kim:2004:MLB

[KL04] Mi-Young Kim and Jong-Hyeok Lee. Machine learning based approach to Sclause segmentation. International Journal of Computer Processing of Oriental Languages (IJCPOL), 17(4):203– ??, December 2004. CODEN ???? ISSN 0219-4279.

Kwong:2008:GEI

[KL08] Oi Yee Kwong and Haizhou Li. Guest Editors’ introduction: Special issue: Selected papers from the Sixth SIGHAN Workshop on Chinese Language Processing. International Journal of Computer Processing Of Languages (IJCPOL), 21(2):97– 99, June 2008. CODEN ???? ISSN 2151-3848.

Kim:2000:ILS

[KLC^{+}00] Byeongchang Kim, Jin-Seok Lee, Jeongwon Cha, Geunbae Lee, and Jong-Hyeok Lee. Integrating linguistic and signal knowledge in a morpheme based speech corpus annotation tool. International Journal of Computer Processing of Oriental Languages (IJCPOL), 13(4):291– ??, 2000. CODEN ???? ISSN 0219-4279.

Kim:2004:MAH

[KLL04] Dong-Il Kim, Jin-Ji Li, and Jong-Hyeok Lee. Matching abnormality in hybrid machine translation. International Journal of Computer Processing of Oriental Languages (IJCPOL), 17(4):253– ??, December 2004. CODEN ???? ISSN 0219-4279.

Kang:2005:TPC

[KNL05] In-Su Kang, Seung-Hoon Na, and Jong-Hyeok Lee. Translation probabilities in cross-language information retrieval. International Journal of Computer Processing of Oriental Languages (IJCPOL), 18(2):115– ??, June 2005. CODEN ???? ISSN 0219-4279.

Kang:2006:CAK

[KNL06] In-Su Kang, Seung-Hoon Na, and Jong-Hyeok Lee. Combination approaches in Korean information retrieval: Words vs. n-grams, and query translation vs. document translation. International Journal of Computer Processing of Oriental Languages (IJCPOL), 19(2–3):153–187, June/September 2006. CO-DEN ???? ISSN 0219-4279.

Kang:2003:IPN

[KOK03] In-Ho Kang, Eun-Jung Oh, and Gil Chang Kim. Incremental PageRanking for newly crawled Web pages. International Journal of Computer Processing of Oriental Languages (IJCPOL), 16(1): 87–??, March 2003. CODEN ???? ISSN 0219-4279.

Keh:2000:EMN

[KSH00] Huan-Chao Keh, Timothy K. Shih, and Jason C. Hung. An electronic mobile notebook on the WWW. International Journal of Computer Processing of Oriental Languages (IJCPOL), 13(3):251– ??, 2000. CODEN ???? ISSN 0219-4279.

Kuo:2011:ICS

[Kuo11] June-Jei Kuo. An intelligent Chinese short message input system using one-character-one-key model. International Journal of Computer Processing Of Languages (IJCPOL), 23(2):131–146, June 2011. CODEN ???? ISSN 2151-3848.

Kwok:2001:ELC

[Kwo01] K. L. Kwok. Exploiting the LDC Chinese–English bilingual wordlist for cross language information retrieval. International Journal of Computer Processing of Oriental Languages (IJCPOL), 14(2): 173–191, 2001. CODEN ???? ISSN 0219-4279.

Kwon:2003:HCB

[Kwo03] Oh-Woog Kwon. Hierarchical categorization based on knearest neighbor approach for Web site classification. International Journal of Computer Processing of Oriental Languages (IJCPOL), 16(1): 63–??, March 2003. CODEN ???? ISSN 0219-4279.

Kotani:2009:ACU

[KY09] Katsunori Kotani and Takehiko Yoshimi. Assessing classification using word-alignment distribution for automatic evaluation of machine translation. International Journal of Computer Processing Of Languages (IJCPOL), 22(2– 3):219–236, June/September 2009. CODEN ???? ISSN 2151-3848.

Kim:2003:RST

[KZCJLL03] Dong-Il Kim, Zheng-Cui, Jinji-Li, and Jong-Hyeok Lee. Resolving structural transfer ambiguity in Chinese-to-Korean machine translation. International Journal of Computer Processing of Oriental Languages (IJCPOL), 16(2): 119–??, June 2003. CODEN ???? ISSN 0219-4279.

Lu:2008:ACT

[LAM08] Jia Lu, Masayuki Asahara, and Yuji Matsumoto. Annotation and classification of three-character Chinese synthetic words. International Journal of Computer Processing Of Languages (IJCPOL), 21(2):101–122, June 2008. CODEN ???? ISSN 2151-3848.

Lin:2012:SOD

[LC12] Bai Lin and Ng Bee Chin. Self-other dimension of Chinese shame words. International Journal of Computer Processing Of Languages (IJCPOL), 24(1):51– ??, March 2012. CODEN ???? ISSN 2151-3848.

Li:2004:FMC

[LD04] Yaodong Li and Ruwei Dai. Framework for a mancomputer cooperative information retrieval system for online discussion systems. International Journal of Computer Processing of Oriental Languages (IJCPOL), 17(4): 273–??, December 2004. CO-DEN ???? ISSN 0219-4279.

Li:2011:ILE

[LDZ11] Hanjing Li, Runzhi Dong, and Tiejun Zhao. Inferring the location of an event in large corpus. International Journal of Computer Processing Of Languages (IJCPOL), 23(3):255– ??, September 2011. CODEN ???? ISSN 2151-3848.

Lee:2002:NNA

[Lee02] Yue-Shi Lee. Neural network approach to adaptive learning with an application to Chinese homophone disambiguation. International Journal of Computer Processing of Oriental Languages (IJCPOL), 15(3): 245–??, September 2002. CO-DEN ???? ISSN 0219-4279.

Lee:2004:GEI

[Lee04] Jong-Hyeok Lee. Guest Editor’s introduction: Special issue: Best papers from the 5th China–Korea Joint Symposium on Oriental Language Processing and Pattern Recognition. International Journal of Computer Processing of Oriental Languages (IJCPOL), 17(4):201– ??, December 2004. CODEN ???? ISSN 0219-4279.

Liu:2008:CSS

[LH08] Meichun Liu and Chiayin Hu.

Conceptual schema as semantic link: A frame-based study of Mandarin cognition verbs. International Journal of Computer Processing Of Languages (IJCPOL), 21 (1):55–75, March 2008. CO-DEN ???? ISSN 2151-3848.

Li:2000:LTA

[LHM^{+}00] H. F. Li, N. Heo, K. Moon, J. H. Lee, and G. Lee. Lexical transfer ambiguity resolution using automatically-extracted concept co-occurrence information. International Journal of Computer Processing of Oriental Languages (IJCPOL), 13(1):53–??, 2000. CODEN ???? ISSN 0219-4279.

Li:2000:IMT

[Li00]  Sheng-Tun Li.  On the

Internet-mediated TV-like distance learning services. International Journal of Computer Processing of Oriental Languages (IJCPOL), 13(3): 235–??, 2000. CODEN ???? ISSN 0219-4279.

Li:2008:GEI

[Li08] Wenjie Li. Guest Editor’s introduction: Special issue: Opinion mining, retrieval and extraction. International Journal of Computer Processing Of Languages (IJCPOL), 21(4):293, December 2008. CODEN ???? ISSN 2151-3848.

Li:2009:GEI

[Li09] Wenjie Li. Guest Editor’s introduction: Special issue: Selected papers from the 22nd International Conference on the Computer Processing of Oriental Languages (ICCPOL 2009). International Journal of Computer Processing Of Languages (IJCPOL), 22(2– 3):131–132, June/September 2009. CODEN ???? ISSN 2151-3848.

Long:2002:ICS

[LJAL02] D. Y. Long, W. J. Jia, P. O. Au, and M. Li. Improving Chinese storing text retrieval systems’ security via a novel maximal prefix coding. International Journal of Computer Processing of Oriental Languages (IJCPOL), 15(4): 379–??, 2002. CODEN ???? ISSN 0219-4279.

Lee:2011:GBD

[LJL11] Yong-Hun Lee, Meixun Jin, and Jong-Hyeok Lee. Graphbased dependency parsing using dynamic features in Korean. International Journal of Computer Processing Of Languages (IJCPOL), 23(2):185– 199, June 2011. CODEN ???? ISSN 2151-3848.

Lam:2000:EDD

[LKCK00] H. C. Lam, W. W. Ki, A. L. S.

Chung, and P. Y. Ko. Experience in designing databases for learning Chinese characters. International Journal of Computer Processing of Oriental Languages (IJCPOL), 13(4):351–??, 2000. CODEN ???? ISSN 0219-4279.

Li:2001:NLG

[LKL01] Hui-Feng Li, Sin-Jae Kang, and Jong-Hyeok Lee. Natural language generation in an MT system based on a multilevel processing strategy. International Journal of Computer Processing of Oriental Languages (IJCPOL), 14(1): 29–46, 2001. CODEN ???? ISSN 0219-4279.

Li:2008:MLT

[LKSL08] Haizhou Li, Jin-Shea Kuo, Jian Su, and Chih-Lung Lin. Mining live transliterations using incremental learning algorithms. International Journal of Computer Processing Of Languages (IJCPOL), 21 (2):183–203, June 2008. CO-DEN ???? ISSN 2151-3848.

Lee:2001:GEI

[LL01] Dik Lun Lee and Jong-Hyeok Lee. Guest Editors’ introduction: Special issue: Information retrieval with Asian languages. International Journal of Computer Processing of Oriental Languages (IJCPOL), 14(2):81–??, 2001. CODEN ???? ISSN 0219-4279.

Lee:1996:MLL

[LP96] Seong-Whan Lee and Hee-Seon Park. Multi-lingual large-set Oriental character recognition using a hierarchical neural network classifier. Computer Processing of Chinese and Oriental Languages: an international journal of the Chinese Language Computer Society, 10(2):129– 45, 1996. CODEN CPCLE6. ISSN 0715-9048.

Lee:2002:INL

[LP02] Hodong Lee and Jong C. Park. Interpretation of natural language queries for relational database access with combinatory categorial grammar. International Journal of Computer Processing of Oriental Languages (IJCPOL), 15(3): 281–??, September 2002. CO-DEN ???? ISSN 0219-4279.

Li:2007:CED

[LQLY07] Wenjie Li, Donglei Qian, Qin Lu, and Chunfa Yuan. Chinese entity detection and tracking: The experience in ACE. International Journal of Computer Processing of Oriental Languages (IJCPOL), 20(4):219–236, December 2007. CODEN ???? ISSN 0219-4279.

Li:2005:CAF

[LRKL05] Jin-Ji Li, Ji-Eun Roh, Dong-Il Kim, and Jong-Hyeok Lee. Contrastive analysis and feature selection for Korean modal expression in Chinese–Korean machine translation system. International Journal of Computer Processing of Oriental Languages (IJCPOL), 18(3):227– ??, September 2005. CODEN ???? ISSN 0219-4279.

Lee:2002:PSB

[LS02a] Sangkon Lee and Masami Shishibori. Passage segmentation based on topic matter. International Journal of Computer Processing of Oriental Languages (IJCPOL), 15(3): 305–??, September 2002. CO-DEN ???? ISSN 0219-4279.

Lee:2002:KSA

[LS02b] Songwook Lee and Jungyun Seo. Korean speech act analysis system using hidden Markov model with decision trees. International Journal of Computer Processing of Oriental Languages (IJCPOL), 15(3):231– ??, September 2002. CODEN ???? ISSN 0219-4279.

Liu:2006:DCV

[LS06] Chang Liu and Zachary Sloane. Developing a core vocabulary for a Mandarin Chinese AAC system using word frequency data. International Journal of Computer Processing of Oriental Languages (IJCPOL), 19(4):285– 300, December 2006. CODEN ???? ISSN 0219-4279.

Li:2008:RVM

[LTB08] Yan-Ping Li, Zhen-Min Tang, and Qian Bo. Research of vowel mapping theory for speaker identification of Chinese Mandarin. International Journal of Computer Processing Of Languages (IJCPOL), 21(3):239– 258, September 2008. CO-DEN ???? ISSN 2151-3848.

Lin:2002:DIC

[LTT02] Yao-Tsung Lin, Shian-Shyong Tseng, and Chang-Jiun Tsai. The design and implementation of a computer-assisted learning expert system. International Journal of Computer Processing of Oriental Languages (IJCPOL), 15(1): 33–??, 2002. CODEN ???? ISSN 0219-4279.

Li:2001:WDF

[LTZ^{+}01] Tong Li, Changjie Tang, Zuo, Jie, et al. Web document filtering technique based on natural language understanding. International Journal of Computer Processing of Oriental Languages (IJCPOL), 14(3): 279–291, 2001. CODEN ???? ISSN 0219-4279.

Lu:2008:GEI

[Lu08] Qin Lu. Guest Editor’s introduction: Special issue: Best papers from the 2007 Chinese Lexical Semantics Workshop (CLSW2007). International Journal of Computer Processing Of Languages (IJCPOL), 21(1):1–2, March 2008. CO-DEN ???? ISSN 2151-3848.

Lunde:1993:HJC

[Lun93] Ken R. Lunde. The history of the Japanese character set and its encoding. Computer Processing of Chinese and Oriental Languages: an international journal of the Chinese Language Computer Society, 7 (1):85–94, June 1993. CO-DEN CPCLE6. ISSN 0715-9048.

Lai:2000:UWP

[LW00] Y. S. Lai and C. H. Wu.

Unknown word and phrase extraction using a phraselike-unit-based likelihood ratio. International Journal of Computer Processing of Oriental Languages (IJCPOL), 13(1):83–??, 2000. CODEN ???? ISSN 0219-4279.

Li:2006:ITD

[LWLW06] Wenjie Li, Mingli Wu, Qin Lu, and Kam-Fai Wong. Integrating temporal distribution information into eventbased summarization. International Journal of Computer Processing of Oriental Languages (IJCPOL), 19(2– 3):201–222, June/September 2006. CODEN ???? ISSN 0219-4279.

Li:2000:ITC

[LWY00] Wenjie Li, Kam-Fai Wong, and Chunfa Yuan. Identifying temporal components in a Chinese temporal information system. International Journal of Computer Processing of Oriental Languages (IJCPOL), 13(2):113– ??, 2000. CODEN ???? ISSN 0219-4279.

Li:2003:DTE

[LWY03] Wenjie Li, Kam-Fai Wong, and Chunfa Yuan. A design of temporal event extraction from Chinese financial news. International Journal of Computer Processing of Oriental Languages (IJCPOL), 16(1): 21–??, March 2003. CODEN ???? ISSN 0219-4279.

Li:2011:EAS

[LWZF11] Binyang Li, Kam-Fai Wong, Lanjun Zhou, and Shi Feng. An efficient approach for sentence-based opinion retrieval. International Journal of Computer Processing Of Languages (IJCPOL), 23(4): 373–??, December 2011. CO-DEN ???? ISSN 2151-3848.

Lee:2003:TSC

[LYK03] Hyun Ah Lee, Juntae Yoon, and Gil Chang Kim. Translation selection by combining multiple measures for sense disambiguation and word selection. International Journal of Computer Processing of Oriental Languages (IJCPOL), 16(3):219– ??, September 2003. CODEN ???? ISSN 0219-4279.

Lin:2001:LVC

[LYY01] Hwei-Jen Lin, Fu-Wen Yang, and Shwu-Huey Yen. Off-line verification for Chinese signatures. International Journal of Computer Processing of Oriental Languages (IJCPOL), 14(1):17–28, 2001. CODEN ???? ISSN 0219-4279.

Lanham:2003:CIO

[LZ03] Elicia Lanham and Wanlei Zhou. Cultural issues in online learning — is blended learning a possible solution? International Journal of Computer Processing of Oriental Languages (IJCPOL), 16(4):

275–??, December 2003. CO-DEN ???? ISSN 0219-4279.

Mohanty:2005:APS

[MBBS05] Sanghamitra Mohanty, Suman Bhattacharya, Sumit Bose, and Sabyasachi Swain. An approach to proper speech segmentation for quality improvement in concatenative text-to-speech system for Indian languages. International Journal of Computer Processing of Oriental Languages (IJCPOL), 18(1):41– ??, March 2005. CODEN ???? ISSN 0219-4279.

Mamish:2009:CAO

[MC09] Safeya Mamish and Mohamed Cheriet. Correcting Arabic OCR errors using improved topic-based language models. International Journal of Computer Processing Of Languages (IJCPOL), 22 (4):321–340, December 2009. CODEN ???? ISSN 2151-3848.

Mochizuki:2000:ERA

[MH00] H. Mochizuki and Y. Hayashi. An efficient retrieval algorithm of compound words using extendable hashing. International Journal of Computer Processing of Oriental Languages (IJCPOL), 13(1): 15–??, 2000. CODEN ???? ISSN 0219-4279.

Mori:2007:ACQ

[MKI07] Tatsunori Mori, Shimpei Kawaguchi, and Madoka Ishioroshi. Answering contextual questions based on the cohesion with knowledge. International Journal of Computer Processing of Oriental Languages (IJCPOL), 20(2– 3):115–135, June/September 2007. CODEN ???? ISSN 0219-4279.

Moon:2002:TDM

[ML02] Kyonghi Moon and Jong-Hyeok Lee. Translation of discontinuous multi-word translation units in a Koreanto-Japanese machine translation system. International Journal of Computer Processing of Oriental Languages (IJCPOL), 15(1):79–??, 2002. CODEN ???? ISSN 0219-4279.

Mizobuchi:2000:CQR

[MMFA00] Shoji Mizobuchi, Kazuhiro Morita, Masao Fuketa, and Jun-Ichi Aoe. Conceptual and quantitative representations of time expressions. International Journal of Computer Processing of Oriental Languages (IJCPOL), 13(4): 313–??, 2000. CODEN ???? ISSN 0219-4279.

Meng:2001:GEI

[MW01] Xiaofeng Meng and Kam Fai Wong. Guest Editors’ introduction: Special issue: Chinese database systems. International Journal of Computer Processing of Oriental Languages (IJCPOL), 14(3):

211–??, 2001. CODEN ???? ISSN 0219-4279.

Meng:2001:OCN

[MWW01] Xiaofeng Meng, Shan Wang, and Kam Fai Wong. Overview of a Chinese natural language interface to databases. International Journal of Computer Processing of Oriental Languages (IJCPOL), 14(3): 213–232, 2001. CODEN ???? ISSN 0219-4279.

Maraoui:2011:UNT

[MZA11] Mohsen Maraoui, Mounir Zrigui, and Georges Antoniadis. Use of NLP tools in CALL system for Arabic. International Journal of Computer Processing Of Languages (IJCPOL), 23(4):427– ??, December 2011. CODEN ???? ISSN 2151-3848.

Maraoui:2012:UNT

[MZA12] Mohsen Maraoui, Mounir Zrigui, and Georges Antoniadis. Use of NLP tools in CALL system for Arabic. International Journal of Computer Processing Of Languages (IJCPOL), 24(2):153– ??, June 2012. CODEN ???? ISSN 2151-3848.

Nguyen:2005:AED

[NH05] Minh Le Nguyen and Susumu Horiguchi. Accuracy enhancement for the decomposition of human-written summary. International Journal of Computer Processing of Oriental Languages (IJCPOL), 18(1):

53–??, March 2005. CODEN ???? ISSN 0219-4279.

Nakamura:2006:CCS

[NIA06] Kenryu Nakamura, Mamoru Iwabuchi, and Norman Alm. A cross-cultural study on the interpretation of picturebased sentences. International Journal of Computer Processing of Oriental Languages (IJCPOL), 19(4):239– 248, December 2006. CODEN ???? ISSN 0219-4279.

Nguyen:2007:STM

[NSNN07] Thai Phuong Nguyen, Akira Shimazu, Minh Le Nguyen, and Vinh Van Nguyen. A syntactic transformation model for statistical machine translation. International Journal of Computer Processing of Oriental Languages (IJCPOL), 20(2–3):79–99, June/September 2007. CO-DEN ???? ISSN 0219-4279.

Oh:2003:SMA

[OC03] Jong-Hoon Oh and Key-Sun Choi. A statistical model for automatic extraction of Korean transliterated foreign words. International Journal of Computer Processing of Oriental Languages (IJCPOL), 16(1):41– ??, March 2003. CODEN ???? ISSN 0219-4279.

Okadome:2007:PED

[Oka07] Takeshi Okadome. A performance evaluation on DSK and qwerty keyboards. International Journal of Computer Processing of Oriental Languages (IJCPOL), 20(1): 15–35, March 2007. CODEN ???? ISSN 0219-4279.

Obied:2012:CBR

[ONAK12] Hanan N. Abu Obied, Maryam S. Nuser, and Mohammed N. Al-Kabi. A comparison between rule and dictionary based romanization of Arabic names. International Journal of Computer Processing Of Languages (IJCPOL), 24 (2):167–??, June 2012. CO-DEN ???? ISSN 2151-3848.

Park:2002:SDS

[PHO02] Jeong-Sun Park, Ki-Cheon Hong, and Il-Seok Oh. Shape decomposition and stroke extraction of handwritten Korean characters. International Journal of Computer Processing of Oriental Languages (IJCPOL), 15(4):469– ??, 2002. CODEN ???? ISSN 0219-4279.

Pong:1995:HCF

[PLYZ95] Man-Chi Pong, Fung Fung Lee, R. Yeung, and Yongguang Zhang. Han character font sharing across incompatible bitmap file formats. Computer Processing of Chinese and Oriental Languages: an international journal of the Chinese Language Computer Society, 9(2):145–153, December 1995. CODEN CPCLE6. ISSN 0715-9048.

Pisarn:2005:TSR

[PTCC05] Chutima Pisarn, Thanaruk Theeramunkong, Nick Cercone, and Junalux Chalidabhongse. Thai spelling recognition using a continuous speech corpus. International Journal of Computer Processing of Oriental Languages (IJCPOL), 18(4):243–??, December 2005. CODEN ???? ISSN 0219-4279.

Peng:2002:XBD

[PWK02] Xu Peng, Yang Wenjun, and Wang Kehong. XMLbased data rendering engine for Web and mobile application scenarios. International Journal of Computer Processing of Oriental Languages (IJCPOL), 15(4):433– ??, 2002. CODEN ???? ISSN 0219-4279.

Qu:2008:WMC

[QLZ^{+}08] Weiguang Qu, Sujian Li, Xiaojun Zhang, Yunfang Wu, Zhifang Sui, and Shiwen Yu. A WSD model for corpus construction. International Journal of Computer Processing Of Languages (IJCPOL), 21 (1):43–53, March 2008. CO-DEN ???? ISSN 2151-3848.

Ramaraj:2005:IAS

[RC05] E. Ramaraj and E. Chandra. Influence of acoustics in speech recognition for oriental languages. International Journal of Computer Processing of Oriental Languages (IJCPOL), 18(4):265– ??, December 2005. CODEN ???? ISSN 0219-4279.

Roh:2006:ECR

[RL06] Ji-Eun Roh and Jong-Hyeok Lee. Examining Cf-ranking methods for text structuring and pronominalization in Korean. International Journal of Computer Processing of Oriental Languages (IJCPOL), 19(1):39– ??, March 2006. CODEN ???? ISSN 0219-4279. URL http:

                                        //ejournals.wspc.com.sg/
                                        ijcpol/19/1901/S0219427906001402.
                                        html.

Ren:2008:EAC

[RZWX08] Feiliang Ren, Jingbo Zhu, Huizhen Wang, and Tong Xiao. An effective approach for coreference resolution. International Journal of Computer Processing Of Languages (IJCPOL), 21(2):135– 149, June 2008. CODEN ???? ISSN 2151-3848.

Saha:2005:EBT

[Sah05] Goutam Kumar Saha. English to Bangla translator: The BANGANUBAD. International Journal of Computer Processing of Oriental Languages (IJCPOL), 18(4):281– ??, December 2005. CODEN ???? ISSN 0219-4279.

Sakai:2001:JEC

[Sak01] Tetsuya Sakai. Japanese-English cross-language information retrieval using machine translation and pseudorelevance feedback. International Journal of Computer Processing of Oriental Languages (IJCPOL), 14(2):83– 107, 2001. CODEN ???? ISSN 0219-4279.

Sakai:2005:ATI

[Sak05] Tetsuya Sakai. Advanced technologies for information access. International Journal of Computer Processing of Oriental Languages (IJCPOL), 18(2):95–??, June 2005. CODEN ???? ISSN 0219-4279.

Sarhan:2009:ACR

[Sar09] Ahmad M. Sarhan. Arabic character recognition using a combination of k-means and k-NN algorithms. International Journal of Computer Processing Of Languages (IJCPOL), 22(4):305– 320, December 2009. CODEN ???? ISSN 2151-3848.

Suen:2005:AIO

[SBN^{+}05] C. Y. Suen, S. Bergler, N. Nobile, W. Pan, and B. Waked. Automatic identification of oriental and other scripts in image documents. International Journal of Computer Processing of Oriental Languages (IJCPOL), 18(2):77– ??, June 2005. CODEN ???? ISSN 0219-4279.

Su:2011:RTR

[SCH11] Qi Su, Helen Kai-Yun Chen, and Chu-Ren Huang. A research on the text reliability based on evidentiality. International Journal of Computer Processing Of Languages (IJCPOL), 23(2):201– 214, June 2011. CODEN ???? ISSN 2151-3848.

Simon:2008:TNE

[ŠHHH08] Petr Šimon, Chu-Ren Huang, Shu-Kai Hsieh, and Jia-Fei Hong. Transliterated named entity recognition based on Chinese word sketch. International Journal of Computer Processing Of Languages (IJCPOL), 21(1):19– 30, March 2008. CODEN ???? ISSN 2151-3848.

Song:2005:TSB

[SJM05] Sa-Kwang Song, Dong Hyun Jang, and Sung Hyon Myaeng. Text summarization based on sentence clustering with rhetorical structure information. International Journal of Computer Processing of Oriental Languages (IJCPOL), 18(2):153–??, June 2005. CO-DEN ???? ISSN 0219-4279.

Shih:2000:QAC

[SKW^{+}00] Timothy K. Shih, Huan-Chao Keh, Ching-Sheng Wang, Anthony Y. Chang, and Chuan-Ho Kao. A qualitative approach to content-based color image retrieval with Chinese caption. International Journal of Computer Processing of Oriental Languages (IJCPOL), 13(2):131– ??, 2000. CODEN ???? ISSN 0219-4279.

Sheard:2003:ALD

[SL03] Judithe Sheard and Julianne Lynch. Accommodating learner diversity in Web-based learning environments: Imperatives for future developments. International Journal of Computer Processing of Oriental Languages (IJCPOL), 16(4):243–??, December 2003. CODEN ???? ISSN 0219-4279.

Shih:2000:GEI

[SM00] Timothy K. Shih and Jianhua Ma. Guest Editors’ introduction: Special issue on virtual university. International Journal of Computer Processing of Oriental Languages (IJCPOL), 13(3):203– ??, 2000. CODEN ???? ISSN 0219-4279.

Shih:2002:GEI

[SM02] Timothy Shih and Jianhua Ma. Guest Editors’ introduction: Special issue: Web and WAP oriental languages multimedia computing. International Journal of Computer Processing of Oriental Languages (IJCPOL), 15(4): 359–??, 2002. CODEN ???? ISSN 0219-4279.

Siqueira:2004:ISL

[SMB04] Sean W. M. Siqueira, Rubens N.

Melo, and Maria Helena L. B. Braz. Increasing the semantics of learning objects. International Journal of Computer Processing of Oriental Languages (IJCPOL), 17(1): 27–??, March 2004. CODEN ???? ISSN 0219-4279.

Selvam:2009:SPN

[SNT09] M. Selvam, A. M. Natarajan, and R. Thangarajan. Structural parsing of natural language text in Tamil language using dependency model. International Journal of Computer Processing Of Languages (IJCPOL), 22(2– 3):237–256, June/September 2009. CODEN ???? ISSN 2151-3848.

Shaalan:2004:MTE

[SRMB04] Khaled Shaalan, Ahmed Rafea, Azza Abdel Moneim, and Hoda Baraka. Machine translation of English noun phrases into Arabic. International Journal of Computer Processing of Oriental Languages (IJCPOL), 17(2):121– ??, June 2004. CODEN ???? ISSN 0219-4279.

Sukhahuta:2001:IES

[SS01] Rattasit Sukhahuta and Dan Smith. Information extraction strategies for Thai documents. International Journal of Computer Processing of Oriental Languages (IJCPOL), 14(2): 153–172, 2001. CODEN ???? ISSN 0219-4279.

Sari:2007:SAL

[SS07] Toufik Sari and Mokhtar Sellami. State-of-the-art of off-line Arabic handwriting segmentation. International Journal of Computer Processing of Oriental Languages (IJCPOL), 20(1):53– 73, March 2007. CODEN ???? ISSN 0219-4279.

Si:2011:EGL

[SS11] Xiance Si and Maosong Sun.

Exploring the granularity level of social tags. International Journal of Computer Processing Of Languages (IJCPOL), 23(3):273– ??, September 2011. CODEN ???? ISSN 2151-3848.

Shin:2005:HCC

[SSH05] Jungpil Shin, Kazunori Suzuki, and Atsushi Hasegawa. Handwritten Chinese character font generation based on stroke correspondence. International Journal of Computer Processing of Oriental Languages (IJCPOL), 18(3):211– ??, September 2005. CODEN ???? ISSN 0219-4279.

Shaalan:2007:AMG

[STK07] Khaled Shaalan, Habib Talhami, and Ibrahim Kamel. Automatic morphological generation for the indexing of Arabic speech recordings. International Journal of Computer Processing of Oriental Languages (IJCPOL), 20(1): 1–14, March 2007. CODEN ???? ISSN 0219-4279.

Shyuu:2000:UWD

[SW00] Jyh-Shing Shyuu and Jhing-Fa Wang. Using wavelet decomposition in Mandarin speech recognition based on a channel-weighting method. International Journal of Computer Processing of Oriental Languages (IJCPOL), 13(2): 159–??, 2000. CODEN ???? ISSN 0219-4279.

Su:2007:MFB

[SZSY07] Qi Su, Ya Zhu, Bin Swen, and Shiwen Yu. Mining featurebased opinion expressions by mutual information approach. International Journal of Computer Processing of Oriental Languages (IJCPOL), 20(2– 3):137–150, June/September 2007. CODEN ???? ISSN 0219-4279.

Tung:2002:ICI

[TCC02] Chun-Liang Tung, Tung-Shou Chen, and Yung-Chang Chiang. An intellectual Chinese input system for mobile phones based upon Chinese phonetic symbols. International Journal of Computer Processing of Oriental Languages (IJCPOL), 15(4): 395–??, 2002. CODEN ???? ISSN 0219-4279.

Tang:2012:TIF

[TH12] Kuo-Ming Tang and Shu-Kai Hsieh. Toward an integrated framework of lexicalized ontology for Buddhist digital archives. International Journal of Computer Processing Of Languages (IJCPOL), 24 (1):103–??, March 2012. CO-DEN ???? ISSN 2151-3848.

Tsou:2005:SCA

[TKWL05] Benjamin K. Y. Tsou, Oi Yee Kwong, Wei Lung Wong, and Tom B. Y. Lai. Sentiment and content analysis of Chinese news coverage. International Journal of Computer Processing of Oriental Languages (IJCPOL), 18(2):171– ??, June 2005. CODEN ???? ISSN 0219-4279.

Tsou:2006:CST

[TLSC06] Benjamin K. Tsou, Tom B. Y.

Lai, K. K. Sin, and Lawrence Y. L. Cheung. Court stenography-to-text (“STT”) in Hong Kong: A jurilinguistic engineering effort. International Journal of Computer Processing of Oriental Languages (IJCPOL), 19(2– 3):99–107, June/September 2006. CODEN ???? ISSN 0219-4279.

Takeuchi:2000:JOE

[TM00] K. Takeuchi and Y. Matsumoto. Japanese OCR error correction using stochastic morphological analyzer and probabilistic word Ngram model. International Journal of Computer Processing of Oriental Languages (IJCPOL), 13(1):69–??, 2000. CODEN ???? ISSN 0219-4279.

Tsuji:2002:AET

[Tsu02]  Keita Tsuji.  Automatic

extraction of translational Japanese-Katakana and English word pairs. International Journal of Computer Processing of Oriental Languages (IJCPOL), 15(3):261– ??, September 2002. CODEN ???? ISSN 0219-4279.

Tzeng:2000:CCC

[TT00] Chih-Hsuan Tzeng and Wen-Hsiang Tsai. Compression of Chinese character patterns in document images based on rectangular region partitioning using contour information and Huffman coding. International Journal of Computer Processing of Oriental Languages (IJCPOL), 13(2):177– ??, 2000. CODEN ???? ISSN 0219-4279.

Velek:2002:VIT

[VNL02] Ondřej Velek, Masaki Nakagawa, and Cheng-Lin Liu. Vector-to-image transformation of character patterns for on-line and off-line recognition. International Journal of Computer Processing of Oriental Languages (IJCPOL), 15(2):187–??, June 2002. CO-DEN ???? ISSN 0219-4279.

Wakahara:2002:HJC

[Wak02] Toru Wakahara. Handwritten Japanese character recognition using adaptive shape normalization by global affine transformation. International Journal of Computer Processing of Oriental Languages (IJCPOL), 15(2):117– ??, June 2002. CODEN ???? ISSN 0219-4279.

Wang:2004:SIL

[Wan04] Feng-Hsu Wang. On scaffolding inductive learning with individualized inquiry examples by machine-learning tools. International Journal of Computer Processing of Oriental Languages (IJCPOL), 17(1): 9–??, March 2004. CODEN ???? ISSN 0219-4279.

Wang:2001:CBL

[WC01] Hsin-Min Wang and Berlin Chen. Content-based language models for spoken document retrieval. International Journal of Computer Processing of Oriental Languages (IJCPOL), 14(2):193– 209, 2001. CODEN ???? ISSN 0219-4279.

Wang:2000:BCW

[WCC00] H. M. Wang, Y. H. Chou, and B. Chen. Browsing the Chinese Web pages using Mandarin speech. International Journal of Computer Processing of Oriental Languages (IJCPOL), 13(1):35– ??, 2000. CODEN ???? ISSN 0219-4279.

Wong:2001:ITB

[WCC01] Kam-Fai Wong, Timothy Kun-Chung Chan, and Chun-Hung Cheng. An investigation on transformationbased error-driven learning algorithm for Chinese noun phrase extraction. International Journal of Computer Processing of Oriental Languages (IJCPOL), 14(1):47– 69, 2001. CODEN ???? ISSN 0219-4279.

Wang:2011:SRM

[WH11] Shan Wang and Chu-Ren Huang. Sense representation in MARVS: a case study on the polysemy of chī. International Journal of Computer Processing Of Languages (IJCPOL), 23(3):285– ??, September 2011. CODEN ???? ISSN 2151-3848.

Wang:2004:HCA

[WHSN04] Chunheng Wang, Yoshinobu Hotta, Misako Suwa, and Satoshi Naoi. Handwritten Chinese address recognition based on enhanced key character extraction and holistic word matching. International Journal of Computer Processing of Oriental Languages (IJCPOL), 17(4):287– ??, December 2004. CODEN ???? ISSN 0219-4279.

Wang:2008:SVW

[WHZ08] Zhenxing Wang, Changning Huang, and Jingbo Zhu. A study on in-vocabulary word segmentation. International Journal of Computer Processing Of Languages (IJCPOL), 21(2):123–134, June 2008. CODEN ???? ISSN 2151-3848.

Wang:2005:CUW

[WL05] Zhuoran Wang and Ting Liu. Chinese unknown word identification based on local bigram model. International Journal of Computer Processing of Oriental Languages (IJCPOL), 18(3):185– ??, September 2005. CODEN ???? ISSN 0219-4279.

Wang:2011:LUT

[WL11] Xiaolin Wang and Yingwei Luo. Locating unregistered toponym for Web map services. International Journal of Computer Processing Of Languages (IJCPOL), 23(2):169– 184, June 2011. CODEN ???? ISSN 2151-3848.

Wang:2003:WBE

[WLJ03] Jianxin Wang, Weini Lu, and Weijia Jia. A Web-based environment for virtual laboratory with CORBA technology. International Journal of Computer Processing of Oriental Languages (IJCPOL), 16(4): 261–??, December 2003. CO-DEN ???? ISSN 0219-4279.

Wong:2001:WMC

[WLR01] K. F. Wong, B. T. Low, and Yongjie Ren. A workflow model for Chinese business processes. International Journal of Computer Processing of Oriental Languages (IJCPOL), 14(3):233– 258, 2001. CODEN ???? ISSN 0219-4279.

Wang:2003:RID

[WLW03] Chung-Ming Wang, Jiunn-Shyan Lee, and Ren-Jie Wang. On realistic ink diffusion synthesis for a calligraphic learning system. International Journal of Computer Processing of Oriental Languages (IJCPOL), 16(2):105– ??, June 2003. CODEN ???? ISSN 0219-4279.

Wong:2002:TRC

[WLYZ02] Kam-Fai Wong, Wenjie Li, Chunfa Yuan, and Xiaodan Zhu. Temporal representation and classification in Chinese. International Journal of Computer Processing of Oriental Languages (IJCPOL), 15(2): 211–??, June 2002. CODEN ???? ISSN 0219-4279.

Wong:2005:GEI

[Won05] Kam-Fai Wong. Guest Editor’s introduction: Special issue: Advanced technologies for oriental information processing. International Journal of Computer Processing of Oriental Languages (IJCPOL), 18(2):75–??, June 2005. CODEN ???? ISSN 0219-4279.

Wang:2002:MPE

[WT02] Jing-Doo Wang and Jyh-Jong Tsay. Mining periodic events from retrospective Chinese news. International Journal of Computer Processing of Oriental Languages (IJCPOL), 15(4):361–??, 2002. CODEN ???? ISSN 0219-4279.

Wang:2007:CSC

[WWCC07] Rongbo Wang, Xiaohua Wang, Zhiqun Chen, and Zheru Chi. Chunk segmentation of Chinese sentences using a combined statistical and rule-based approach (CSRA). International Journal of Computer Processing of Oriental Languages (IJCPOL), 20(2– 3):197–218, June/September 2007. CODEN ???? ISSN 0219-4279.

Wong:2006:LBS

[WXL06] Kam-Fai Wong, Yunqing Xia, and Wenjie Li. Linguistic and behavioural studies of Chinese Chat language. International Journal of Computer Processing of Oriental Languages (IJCPOL), 19(2– 3):133–152, June/September 2006. CODEN ???? ISSN 0219-4279.

Wong:2005:OTI

[WXLY05] Kam-Fai Wong, Yunqing Xia, Wenjie Li, and Chunfa Yuan. An overview of temporal information extraction. International Journal of Computer Processing of Oriental Languages (IJCPOL), 18(2):137– ??, June 2005. CODEN ???? ISSN 0219-4279.

Wong:2008:PBO

[WXX^{+}08] Kam-Fai Wong, Yunqing Xia, Ruifeng Xu, Mingli Wu, and Wenjie Li. Pattern-based opinion mining for stock market trend prediction. International Journal of Computer Processing Of Languages (IJCPOL), 21(4):347– 361, December 2008. CODEN ???? ISSN 2151-3848.

Xu:2011:GEI

[XK11a] Ruifeng Xu and Chunyu Kit. Guest Editors’ introduction. International Journal of Computer Processing Of Languages (IJCPOL), 23(4):307– ??, December 2011. CODEN ???? ISSN 2151-3848.

Xu:2011:LML

[XK11b] Ruifeng Xu and Chunyu Kit. Learning multiple level features for opinion analysis. International Journal of Computer Processing Of Languages (IJCPOL), 23(4):349– ??, December 2011. CODEN ???? ISSN 2151-3848.

Xu:2009:BCC

[XLWL09] Ruifeng Xu, Qin Lu, Kam-Fai Wong, and Wenjie Li. Building a Chinese collocation bank. International Journal of Computer Processing Of Languages (IJCPOL), 22(1):21– 47, March 2009. CODEN ???? ISSN 2151-3848.

Xia:2008:SVS

[XWW08] Yunqing Xia, Linlin Wang, and Kam-Fai Wong. Sentiment vector space model for lyric-based song sentiment classification. International Journal of Computer Processing Of Languages (IJCPOL), 21(4):309– 330, December 2008. CODEN ???? ISSN 2151-3848.

Yang:2011:DBI

[YCW11] Xiaohong Yang, Qingcai Chen, and Xiaolong Wang. A dictionary based indexing approach for music information retrieval. International Journal of Computer Processing Of Languages (IJCPOL), 23(4):389–??, December 2011. CODEN ???? ISSN 2151-3848.

Yuan:2001:EMC

[YCWL01] Chunfa Yuan, Gang Chen, K. F. Wong, and Wenjie Li. An effective method for Chinese grammar rule learning. International Journal of Computer Processing of Oriental Languages (IJCPOL), 14(4): 361–??, 2001. CODEN ???? ISSN 0219-4279.

Yang:2008:SAW

[YLC08] Changhua Yang, Kevin Hsin-Yih Lin, and Hsin-Hsi Chen. Sentiment analysis in Weblog using contextual information: A machine learning approach. International Journal of Computer Processing Of Languages (IJCPOL), 21 (4):331–345, December 2008. CODEN ???? ISSN 2151-3848.

Yong:2006:COA

[Yon06] Sarah Yong. Comparison of outcomes of an augmentative and alternative communication system used by an English and Mandarin Chinese speaker — A clinical perspective. International Journal of Computer Processing of Oriental Languages (IJCPOL), 19(4):263– 273, December 2006. CODEN ???? ISSN 0219-4279.

Yoon:2002:ESD

[Yoo02]  Juntae Yoon.  Efficient

semi-deterministic parsing for Korean using lexical cooccurrence data from a corpus. International Journal of Computer Processing of Oriental Languages (IJCPOL), 15(4):493–??, 2002. CODEN ???? ISSN 0219-4279.

Yang:2011:GEI

[YY11] Christopher C. Yang and Chin-Sheng Yang. Guest Editors’ introduction. International Journal of Computer Processing Of Languages (IJCPOL), 23(2):107– 108, June 2011. CODEN ???? ISSN 2151-3848.

Yang:2001:DSC

[YYW01] Xiaochun Yang, Ge Yu, and Guoren Wang. Distributed semantic constraints in virtual enterprise information integrating systems. International Journal of Computer Processing of Oriental Languages (IJCPOL), 14(3):259– 278, 2001. CODEN ???? ISSN 0219-4279.

Zhou:2011:GEI

[Zho11] Guodong Zhou. Guest Editor’s introduction. International Journal of Computer Processing Of Languages (IJCPOL), 23(3):215– ??, September 2011. CODEN ???? ISSN 2151-3848.

Zhang:2008:SCU

[ZHW08] Qi Zhang, Xunajing Huang, and Lide Wu. Sentiment classification using supervised and semi-supervised conditional maximum entropy modeling. International Journal of Computer Processing Of Languages (IJCPOL), 21(4):295– 308, December 2008. CODEN ???? ISSN 2151-3848.

Zhou:2011:AIC

[ZLZ11] Yu Zhou, Peng Liu, and Cheng Zong. Approaches to improving corpus quality for statistical machine translation. International Journal of Computer Processing Of Languages (IJCPOL), 23(4):327– ??, December 2011. CODEN ???? ISSN 2151-3848.

Zhang:2005:CJC

[ZMI05] Yujie Zhang, Qing Ma, and Hitoshi Isahara. Construction of a Japanese-Chinese bilingual dictionary using English as an intermediary. International Journal of Computer Processing of Oriental Languages (IJCPOL), 18(1):23– ??, March 2005. CODEN ???? ISSN 0219-4279.

Zhuang:2004:TCE

[ZMWP04] Yueting Zhuang, Yi Mao, Fei Wu, and Yunhe Pan. Towards comprehensive 3D enabled Web-based learning. International Journal of Computer Processing of Oriental Languages (IJCPOL), 17(1): 3–??, March 2004. CODEN ???? ISSN 0219-4279.

Zouaghi:2012:HAA

[ZMZ12] Anis Zouaghi, Laroussi Merhbène, and Mounir Zrigui. A hybrid approach for Arabic word sense disambiguation. International Journal of Computer Processing Of Languages (IJCPOL), 24(2):133– ??, June 2012. CODEN ???? ISSN 2151-3848.

Zhang:2004:QER

[ZN04] Lihua Zhang and Yiu-Kai Ng. A query engine for retrieving information from Chinese HTML documents. International Journal of Computer Processing of Oriental Languages (IJCPOL), 17(3):135– ??, September 2004. CODEN ???? ISSN 0219-4279.

Zhao:2011:AER

[ZQCL11] Yanyan Zhao, Bing Qin, Wanxiang Che, and Ting Liu. Appraisal expression recognition with syntactic path for sentence sentiment classification. International Journal of Computer Processing Of Languages (IJCPOL), 23(1): 21–37, March 2011. CO-DEN ???? ISSN 2151-

3848.  URL http://www.
worldscinet.com/ijcpol/
23/preserved-docs/2301/
S179384061100219X.pdf.

Zhu:2009:LRE

[ZWT09] Jingbo Zhu, Huizhen Wang, and Benjamin K. Tsou. Learning representative examples for data annotation. International Journal of Computer Processing Of Languages (IJCPOL), 22(2– 3):155–169, June/September 2009. CODEN ???? ISSN 2151-3848.

Zhong:2011:WSC

[ZZ11] Ding Zhong and Yangsen Zhang. A word similarity calculation method based on CSD and statistics. International Journal of Computer Processing Of Languages (IJCPOL), 23(3):217– ??, September 2011. CODEN ???? ISSN 2151-3848.

Zhou:2007:BCN

[ZZLZ07] Guodong Zhou, Min Zhang, Junhui Li, and Qiaoming Zhu. Building a collocation net. International Journal of Computer Processing of Oriental Languages (IJCPOL), 20(2– 3):181–196, June/September 2007. CODEN ???? ISSN 0219-4279.

Zhang:2011:ECS

[ZZZL11] Jian Zhang, Hai Zhao, Liqing Zhang, and Bao-Liang Lu. An empirical comparative study on two large-scale hierarchical text classification approaches. International Journal of Computer Processing Of Languages (IJCPOL), 23 (4):309–??, December 2011. CODEN ???? ISSN 2151-3848.